### v4

# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

Abstract
The Lindblad equation embodies a fundamental paradigm of the quantum theory of open systems, and the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) generation theorem says precisely which superoperators can appear on its right-hand side. These are the generators of completely positive trace-preserving (or nonincreasing) semigroups. We prove a generalization, with time-dependent generator, as an application of an investigation of the geometry of the class of completely positive (CP) maps. The treatment of the finite-dimensional setting is based on a basis-free Choi-Jamiołkowski type isomorphism. The infinite-dimensional case is bootstrapped from the finite-dimensional theory via a sequence of finite-dimensional approximations. Kraus decomposition is established along the way, in the guise of an extremal decomposition of the closed convex cone of CP maps. No appeal is made to results from the representation theory of operator algebras.
Lindblad equation, Jamiołkowski isomorphism, complete positivity, Markovian
1 Introduction
The requirement of self-adjointness of the Hamiltonian $H$ in the Schrödinger equation $i\hbar\tfrac{d}{dt}\psi=H\psi$ describing evolution of the state vector of a closed quantum system can be understood as coming directly from the probabilistic interpretation: It is necessary and sufficient in order that the norm of any initial state vector be preserved. The von Neumann equation
$\frac{d\rho}{dt}=\frac{i}{\hbar}[\rho,H],$ (1.1)
follows, given that $\left|{{\psi}}\middle\rangle\middle\langle{{\psi}}\right|$ is the density matrix corresponding to $\psi$, and the demand that convex mixture be respected — a constraint which also has a direct probabilistic interpretation. The Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) generation theorem tells what are the similarly general probabilistic constraints on an “equation of the same sort” for an open system. To explain the quoted term, let us present the von Neumann equation in two parts. The first part is
$\frac{d\rho}{dt}=\mathscr{L}\rho,$ (1.2)
where $\mathscr{L}$ is a linear superoperator, and the second part specifies a particular form for it, namely commutator with an anti-hermitian operator. Now, keep the first part (1.2) and ask what are the minimal constraints on the generator $\mathscr{L}$ for the resulting evolution to make physical sense. An evolution operator $\mathcal{E}$ must preserve positivity of a density operator, but an even more restrictive property, namely complete positivity (CP), is required to ensure that tensor products of operations are coherent in the face of entanglement. We dub the problem of finding necessary and sufficient conditions for a continuous-time CP evolution, the “GKSL problem”.
Five decades ago, Gorini, Kossakowski and Sudarshan[1] gave a solution of the GKSL prolem in finite dimensions, and Lindblad[2] for infinite-dimensional (separable) Hilbert spaces. Like much pioneer work, these are difficult papers, especially that of Lindblad. While the subject of open quantum systems was long the province only of specialists, recent developments have made it relevant to a much broader range of scientists[3, 4, 5, 6, 7]. It is hoped that the present approach may appeal to some of them, as well as to specialists. This approach might be described as geometrical, structural, and basis-free. And, for what concerns the infinite-dimensional case, free of appeals to results in operator-algebraic representation theory. One might go so far as to say that the subject here is really the class of CP operators, particularly its geometry, with the GKSL problem being an application. Not only the endpoint, but also this context, is treated with care. In principle, no prior knowledge of open quantum systems or complete positivity is required of the reader. A somewhat novel introduction to the CP concept is given in §3.2. The finite-dimensional theory is developed in §2 – §5; the remainder of the paper extends it to separable spaces.
Usually, the GKSL theorem is stated thus: A superoperator $\mathscr{L}\colon\mathfrak{B}(\mathfrak{B}(\mathcal{H}))$ generates a quantum dynamical semigroup [trace-preserving CP semigroup] if and only if it can be written in the form
$\mathscr{L}\rho=-[iH,\rho]+\sum_{k}\gamma_{k}\Big\{V_{k}\rho V_{k}^{\dagger}-\tfrac{1}{2}[V_{k}^{\dagger}V_{k},\rho]_{+}\Big\}$ (1.3)
with $\gamma_{k}\geq 0$, $H$ hermitian, and the $V_{k}$ operators on $\mathcal{H}$. Actually, the form on the RHS of (1.3) is what Breuer & Petruccione (p. 122 [3]) call “diagonal form”. Their “first standard form” involves an explicit basis of $\mathfrak{B}(\mathcal{H})$. GKS give only the first standard form. The proof of the “only if” direction in Breuer & Petruccione seems to have become fairly standard. Very similar proofs are found in Rivas & Huelga (pp. 35–37, [4]), Vacchini (pp. 204–207, [5]), Manzano [8], and Alicki & Lendi[9]. This standard proof is a brute-force approach and does not offer much insight, in the author’s opinion. Not all sources treat the “if” direction. GKS rely on a difficult (and somewhat difficult to find) operator-algebraic result [10]. Rivas & Huelga use the Lie-Trotter product formula. (They state the result for time-dependent $\gamma_{k}$ and $V_{k}$ without giving any regularity conditions.) Vacchini gives a clean proof using a Duhamel/Dyson approach.
The approach here is different. A fundamental tool, developed in §2, is a Hilbert space isomorphism $\mathfrak{J}$ which is a basis-free version of the Choi-Jamiołkowski isomorphism. §3 critically introduces the CP concept, which is needed to ensure that we are free to take tensor products of physically admissible quantum operations. The Jamiołkowski isomorphism is specialized to obtain an isomorphism between $\mathsf{CP}(\mathcal{H},\mathcal{K})$, the completely positive maps from $\mathfrak{B}(\mathcal{H})$ to $\mathfrak{B}(\mathcal{K})$, and $\mathrm{Pos}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$, the closed convex pointed cone of positive operators on the Hilbert space $\mathfrak{B}(\mathcal{H},\mathcal{K})$. This requires we are working in finite dimensions. §4 illustrates this point of view with a number of examples. A notable result here is a derivation of Kraus decomposition as extremal decomposition, thus giving meaning to the former.
§5 turns the development toward the GKSL problem. We are interested in CP families of evolution operators $\mathcal{E}(t,s)$ for $t\geq s$ satisfying the differential equation
$\tfrac{\partial}{\partial t}\mathcal{E}(t,s)=\mathscr{L}(t)\mathcal{E}(t,s);\quad\mathcal{E}(s,s)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})};$ (1.4)
and the Markovianity condition $u\geq t\geq s\Rightarrow\mathcal{E}(u,s)=\mathcal{E}(u,t)\circ\mathcal{E}(t,s)$. The problem is to find necessary and sufficient conditions on the generator $\mathscr{L}(t)$. It is easy to see that the generator must at least lie in the tangent cone to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$, denoted $\mathsf{cp}^{+}(\mathcal{H})$. We will show that it is sufficient if $\mathscr{L}(t)$ varies continuously in this tangent cone. Moreover, $\mathsf{cp}^{+}(\mathcal{H})$ coincides with the image (“$L$-cone”) of $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ under the linear map
$L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;],$ (1.5)
and there are linear maps $\Delta$ (“Lindblad parametrizers”) which are right inverse to $L$; and such that the first component of $\Delta\Lambda$ is CP if (and only if) $\Lambda$ is in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ and zero if $\Lambda$ is in the tangent space $\mathsf{cp}(\mathcal{H})$. This latter space consists of the superoperators generating reversible CP evolutions.
Although both Kraus decompositions and Lindblad parametrizations both involve can-be-written type claims, Lindblad parametrizers show that the latter is in a sense much simpler as there are uniform ways of extracting a suitable $(\Psi,G,H)$ parametrization. This is of importance when dealing with time-varying generators or perturbation problems.
The second part of the paper, starting in §7, is devoted to handling the infinite-dimensional (separable) case, bootstrapping from the finite-dimensional theory by way of a sequence of finite-dimensional approximations. After a brief orientation toward the separable setting in §7, the machinery for such approximations is developed in §§8 under the rubric filtration; and 9 derives a metric ($d$-metric) on a base space of a filtration with respect to which the unit ball is compact.
The following two sections demonstrate the utility of that machinery by extending all the basic results to the separable setting. §10 shows that $\mathsf{CP}(\mathcal{H})$ is $d$-closed, that Kraus decompositions exist, and that they are extremal decompositions. It is perhaps surprising that existence of Kraus decomposition can be shown this way since the standard approach[11, 12] makes it seem a fairly deep result of $C^{*}$-algebra representation theory.
Finally, §11 tackles the GKSL problem in the separable setting, avoiding the $C^{*}$-algebraic tools deployed by Lindblad[2]. All the basic results of the finite-dimensional theory are extended: acceptability of generators (norm) continuously varying in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ which coincides with the $L$-operators, and the existence of Lindblad parametrizers. In addition, we show that $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ are $d$-closed (hence, e.g. weak-operator closed).
A table of notations and terminology can be found in Appendix A.
2 Generalized Jamiołkowski isomorphism: the power of swap
This Section develops a version of the Jamiołkowski isomorphism $\mathfrak{J}$ partially inspired by [13]. Although the demonstration uses tensor products, the formulation is entirely in the relevant function spaces, without the tensor product proxies used in references [13] and [14]. Only in the next section will the usefulness of $\mathfrak{J}$ become clear. The key properties (2.9) – (2.11) encapsulate what is really essential for the sequel.
2.1 Notational conventions
Hilbert spaces are denoted by calligraphic letters $\mathcal{H}$, $\mathcal{K}$, etc. Vectors may be denoted by lower-case greek letters, $\psi$$\phi$, and so on, or by lower-case latin letters, matching the Hilbert space name, e.g., $h,h^{\prime}\in\mathcal{H}$, as a helpful mnemonic. Upper-case italic letters $A$, $B$, etc. denote operators. The space of linear operators from $\mathcal{H}$ to $\mathcal{K}$ is mostly denoted $\mathfrak{B}(\mathcal{H},\mathcal{K})$ ($\mathfrak{B}(\mathcal{H},\mathcal{H})$ is abbreviated to $\mathfrak{B}(\mathcal{H})$), but in this section we use $\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)$. Running the arrow right-to-left is somewhat unconventional, but corresponds with the order in which operators are composed and can be helpful when there are many spaces to keep track of. Upper-case greek letters $\Lambda$, $\Gamma$, stand for operators on operators, that is, elements of a space like $\left({\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)}\leftarrow{\left({\mathcal{N}}\leftarrow{\mathcal{M}}\right)}\right)$. As is customary, we usually refer to these as superoperators.
Simple juxtaposition indicates composition, as in $AB$, or application, as in $A\psi$. (Considering $\psi\in\mathcal{H}$ as a linear map $\mathbb{C}$$\rightarrow\mathcal{H}$ taking $1$ to $\psi$, this is also composition.) $\Lambda\cdot\psi\overline{\phi}$ is the same as $\Lambda(\psi\overline{\phi})$; the is a form of punctuation, equivalent to an opening parenthesis with an invisible partner as far to the right as possible. The composition symbol, $\circ$”, will often be omitted unless it is necessary for disambiguation.
[…]
2.2 Spaces of operators
2.2.1 Dual space
$\mathcal{H}^{*}$ denotes the dual space of $\mathcal{H}$, i.e., the space of $\mathbb{C}$}-linear maps $\mathcal{H}\rightarrow{\mathbb{C}}$. This only involves the linear structure of $\mathcal{H}$, not its Hilbert space structure. However, there is a natural Hilbert space structure induced on $\mathcal{H}^{*}$ via a natural conjugate-linear bijection between $\mathcal{H}$ and $\mathcal{H}^{*}$. Dirac writes the latter map as $\left|{\psi}\right\rangle\mapsto\left\langle{\psi}\right|$. Instead, we will use a simple overbar: $\overline{\psi}$ is defined by $\overline{\psi}(\phi)=\left\langle{\psi}\middle|{\phi}\right\rangle$. The parentheses indicate function application, but we dispense with those, too, and write simply $\overline{\psi}{\phi}$. Recall that conjugate-linear means that, for $c\in{\mathbb{C}}$, $\overline{(c\psi)}=\overline{c}\overline{\psi}$. With this notation, the inner-product on $\mathcal{H}^{*}$ is
$\left\langle{\overline{\psi}}\middle|{\overline{\phi}}\right\rangle=\left\langle{{\phi}}\middle|{{\psi}}\right\rangle.$ (2.1)
Both sides are linear in $\psi$ and conjugate-linear in $\phi$.
2.2.2 Hilbert-Schmidt space
$\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)$ has a Hilbert space structure given by the Hilbert-Schmidt inner product
$\left\langle{A}\middle|{B}\right\rangle{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\operatorname{\mathrm{Tr}}A^{\dagger}B,$ (2.2)
where the hermitian adjoint of $A\colon{\mathcal{K}}\leftarrow{\mathcal{H}}$ is the operator $A^{\dagger}\in\left({\mathcal{H}}\leftarrow{\mathcal{K}}\right)$ defined by $\left\langle{A^{\dagger}k}\middle|{h}\right\rangle_{\mathcal{H}}=\left\langle{k}\middle|{Ah}\right\rangle_{\mathcal{K}}$. Subscripting of inner products like this is unnecessary, hence will not be used. For instance, from the type of $A$ in the above, one deduces $h\in\mathcal{H}$ and $k\in\mathcal{K}$. The inner products on such operator spaces satisfy the identities
$\left\langle{A}\middle|{BC}\right\rangle=\left\langle{AC^{\dagger}}\middle|{B}\right\rangle=\left\langle{B^{\dagger}A}\middle|{C}\right\rangle$ (2.3)
for $A\colon{\mathcal{K}}\leftarrow{\mathcal{H}}$, $B\colon{\mathcal{K}}\leftarrow{\mathcal{N}}$, $C\colon{\mathcal{N}}\leftarrow{\mathcal{H}}$. In general, each of the three inner products in the above display are in different spaces.
Matrix elements of an operator $A\colon\mathcal{H}\rightarrow\mathcal{K}$ can be expressed in terms of the associated Hilbert-Schmidt inner product as
$\left\langle{k}\middle|{Ah}\right\rangle=\left\langle{k\overline{h}}\middle|{A}\right\rangle.$ (2.4)
Note that the two inner products here are (LHS) in $\mathcal{K}$ and (RHS) in $\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)$. The maneuver (2.4) will be useful in the sequel.
$\mathfrak{B}(()\mathcal{H},\mathcal{K})$ and $\mathfrak{B}(()\mathcal{M},\mathcal{N})$ thus being themselves Hilbert spaces in a natural way, we can repeat the process ad infinitum. However, an isomorphism is useful for dealing with the inner products here.
2.3 Jamiołkowski transformation
2.3.1 A fundamental isomorphism: $\mathfrak{B}(()\mathcal{H},\mathcal{K})\cong\mathcal{K}\otimes\mathcal{H}^{*}$
There is a natural Hilbert isomorphism, i.e., unitary mapping
$\begin{array}[]{ccc}\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)&\cong&\mathcal{K}\otimes\mathcal{H}^{*}\\ k\overline{h}&\leftrightarrow&k\otimes\overline{h}.\end{array}$ (2.5)
The indicated correspondence between rank-one operators $k\overline{h}$ in $\mathfrak{B}(\mathcal{H},\mathcal{K})$ and decomposable tensors $k\otimes\overline{h}$ in $\mathcal{K}\otimes\mathcal{H}^{*}$ extends to sums of such things linearly. Coherence of the definition is guaranteed by the identical distributivity and scalar multiplication properties: $k\otimes\overline{h}+k^{\prime}\otimes\overline{h}=(k+k^{\prime})\otimes\overline{h}$ and $c(k\otimes\overline{h})=(ck)\otimes\overline{h}=k\otimes\overline{\overline{c}h}$ are still correct equations if $\otimes$’ is replaced by $\circ$’. Thus, (2.5) defines a linear isomorphism. Because it preserves inner products, it is even a Hilbert isomorphism (unitary mapping). To see this, it suffices to verify inner products on dyadics.
[…]
Recall that the inner product on the Hilbert tensor product $\mathcal{M}\otimes\mathcal{N}$ is defined by $\left\langle{m\otimes n}\middle|{m^{\prime}\otimes n^{\prime}}\right\rangle=\left\langle{m}\middle|{m^{\prime}}\right\rangle\left\langle{n}\middle|{n^{\prime}}\right\rangle$.
It may be tempting to consider $k\overline{h}$ and $k\otimes\overline{h}$ as literally identical. That is, however, a serious semantic mistake. $k\overline{h}$’ represents a combination in series, whereas $k\otimes\overline{h}$’ represents a combination in parallel.
2.3.2 Jamiołkowski transformation defined
Now for the point of the tensor products. Swapping the order of factors in a tensor product is trivially a Hilbert space isomorphism. The Jamiołkowski isomorphism
[…]
is obtained by this simple maneuver, together with the isomorphism (2.5). The following display shows in detail how $\mathfrak{J}$ is constructed from elementary isomorphisms, tracking a generic separable element in the middle column, with type shown on the right.
[…]
From the penultimate to final line, the first three steps are simply reversed, and the isomorphism is extended to the entire spaces by linearity.
2.4 Key properties of $\mathfrak{J}$
$\mathfrak{J}\text{ is its own inverse}$ (2.9)
$A\overline{B}\;\xleftrightarrow{\;\mathfrak{J}\;}\;A\,\square\,B^{\dagger}$ (2.10)
$\mathfrak{J}(\Lambda\otimes\Gamma)=\mathfrak{J}\Lambda\otimes\mathfrak{J}\Gamma$ (2.11)
Here is the definition of the new notation.
Definition 2.1.
$S\,\square\,T$ is the superoperator which acts on $X$ to yield the composite
${\mathcal{N}}\xleftarrow{S}{\mathcal{M}}\xleftarrow{X}{\mathcal{K}}\xleftarrow{T}{\mathcal{H}}.$ (2.12)
Thus, with $S$ and $T$ having types as indicated in the diagram,
$S\,\square\,T\colon\left({\mathcal{M}}\leftarrow{\mathcal{K}}\right)\leftarrow\left({\mathcal{N}}\leftarrow{\mathcal{H}}\right).$ (2.13)
The point of this notation is that it allows us to avoid dummy arguments. The mnemonic is that “the box shows where the argument of $S\,\square\,T$ goes”. $\dashv$
Proof of (2.9).
Immediate from the definition. ∎
Proof of (2.10).
Consider the special case
$A=n\overline{k},\quad B=m\overline{h},\quad X=k^{\prime}\overline{h^{\prime}}.$ (2.14)
Then, $\mathfrak{J}(A\overline{B})=\mathfrak{J}(n\overline{k}\,(\overline{m\overline{h}}))=n\overline{m}\cdot\overline{k\overline{h}}$, so
$\begin{split}\mathfrak{J}(A\overline{B})X=\,&\Big(n\overline{m}\,(\overline{k\overline{h}})\Big)k^{\prime}\overline{h^{\prime}}\\ =\,&n\overline{m}\left\langle{k\overline{h}}\middle|{k^{\prime}\overline{h^{\prime}}}\right\rangle.\end{split}$ (2.15)
On the other hand,
$\begin{split}n\overline{k}(k^{\prime}\overline{h^{\prime}}){h\overline{m}}&=n\left\langle{k}\middle|{k^{\prime}}\right\rangle\left\langle{h^{\prime}}\middle|{h}\right\rangle\overline{m}\\ &=n\overline{m}\left\langle{k\overline{h}}\middle|{k^{\prime}\overline{h^{\prime}}}\right\rangle.\end{split}$ (2.16)
Hence, since $(m\overline{h})^{\dagger}=h\overline{m}$, the special case (2.14) of $\mathfrak{J}(A\overline{B})X=AXB^{\dagger}$ is proven. But both sides of this special case equation are linear in $A$ and $X$ and conjugate linear in $B$, so this is enough. ∎
Proof of (2.11).
By (2.10),
$A\overline{B}\otimes S\overline{T}=(A\otimes S)(\overline{B\otimes T})\;\xleftrightarrow{\;\mathfrak{J}\;}\;(A\otimes S)\,\square\,(B^{\dagger}\otimes T^{\dagger}).$ (2.17)
On the other hand,
$\begin{split}(A\otimes S)(X\otimes Y)(B^{\dagger}\otimes T^{\dagger})=\,&AXB^{\dagger}\otimes SYT^{\dagger}\\ =\,&\mathfrak{J}(A\overline{B})X\otimes\mathfrak{J}(S\overline{T})Y\\ =\,&(\mathfrak{J}(A\overline{B})\otimes\mathfrak{J}(S\overline{T}))(X\otimes Y).\end{split}$ (2.18)
Thus, $\mathfrak{J}(A\overline{B}\otimes S\overline{T})=\mathfrak{J}(A\overline{B})\otimes\mathfrak{J}(S\overline{T})$. This suffices by linearity ($\Lambda\otimes\Gamma$ is a sum of terms like those). ∎
Since any superoperator of type $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}),\mathfrak{B}(\mathcal{M},\mathcal{N}))$ can be written as a sum of superoperators $A\overline{B}$ for $A\colon\mathfrak{B}(\mathcal{M},\mathcal{N})$ and $B\colon\mathfrak{B}(\mathcal{H},\mathcal{K})$, (2.10) is equivalent to the definition of $\mathfrak{J}$. Now that we have (2.10), we will not need to make direct appeal to the definition anymore.
Example 2.1 (Orthonormal bases).
Here is a simple application of $\mathfrak{J}$; it will not be used in the main line of development. $\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)$ is spanned by rank-one operators, i.e., those of the form $k\overline{h}$. From these, according to (2.6), we can extract an orthonormal basis (ONB) $\{k_{i}\overline{h_{j}}\}$, where $\{k_{i}\}$ and $\{h_{i}\}$ are orthonormal bases of $\mathcal{K}$ and $\mathcal{H}$, respectively.
Now, specialize this to the case of spaces of form $\mathfrak{B}(\mathcal{H},\mathcal{K})$ and $\mathfrak{B}(\mathcal{M},\mathcal{N})$. If $\{A_{i}\}$ is an ONB of $\mathfrak{B}(\mathcal{H},\mathcal{K})$, and $\{S_{\alpha}\}$ of $\mathfrak{B}(\mathcal{M},\mathcal{N})$, then $\{A_{i}\overline{S_{\alpha}}\}$ is an ONB of $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}),\mathfrak{B}(\mathcal{M},\mathcal{N}))$. Because $\mathfrak{J}$ is a Hilbert isomorphism, $\{A_{i}\,\square\,S_{\alpha}^{\dagger}\}$, is also an ONB. $\dashv$
2.5 Operational significance of the Hilbert-Schmidt inner product
We make a short aside here on the physical/operational significance of the inner products we are working with. Here and throughout the paper, we will be using the following notation.
Definition 2.2 ($\bm{\pi}$}).
For $h$ in Hilbert space $\mathcal{H}$,
${\bm{\pi}}[h]{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}h\overline{h}$ (2.19)
is a member of $\mathfrak{B}(\mathcal{H})$ proportional to the orthoprojector onto $h$, but not equal to it unless $h$ is normalized. $\dashv$
In quantum mechanics, the inner product on a state-vector-space $\mathcal{H}$ is fundamental, and the Born interpretation explains its physical significance: $|\left\langle{\phi}\middle|{\psi}\right\rangle|^{2}$ is a transition probability. If we prepare a projective measurement to test whether the state is ${\bm{\pi}}(\phi)$, the probability of success, assuming the state is actually ${\bm{\pi}}(\psi)$ is $|\left\langle{\phi}\middle|{\psi}\right\rangle|$. But, this is just $\left\langle{{\bm{\pi}}(\phi)}\middle|{{\bm{\pi}}{\psi}}\right\rangle$ (no modulus squared). It looks as if the Hilbert-Schmidt inner product on $\mathfrak{B}(\mathcal{H})$ (or at least $\mathrm{Pos}(\mathcal{H})$) might be at least as well-founded as that on $\mathcal{H}$. But what of the extension to general (i.e., mixed) states? Consider such a normalized state $\rho=\sum c_{i}{\bm{\pi}}(\psi_{i})$, with normalized $\psi$. Preparing this state is generally accepted as equivalent to preparing ${\bm{\pi}}(\psi_{i})$ with probability $c_{i}$. In that case, $\left\langle{{\bm{\pi}}(\phi)}\middle|{\rho}\right\rangle$ still makes sense as a success probability. To replace $\phi$ with a mixed state requires a little asymmetry. We should interpret $\sigma=\sum d_{i}{\bm{\pi}}(\phi_{i})$ as a randomized measurment. Then, $\left\langle{\sigma}\middle|{\rho}\right\rangle$ is again the corresponding success probability.
The operational credentials of the induced inner product on $\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ are much murkier. The best we can do is to observe that $\left\langle{A\overline{B}}\middle|{\Lambda}\right\rangle=\left\langle{A}\middle|{\Lambda B}\right\rangle$, where the inner product on the right is already interpreted. Perhaps this failure is to be expected, however. While the measurment probability interpretation continuous to work in an infinite-dimensional context, the Hilbert space in $\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ is too small to be useful.
3 Jamiołkowski correspondence between classes of operators
In developing the generalized Jamiołkowski isomorphism, we used four different underlying Hilbert spaces ($\mathcal{H}$, $\mathcal{K}$, $\mathcal{M}$, $\mathcal{N}$) because that helped to keep the distinct rôles of the spaces clearer, and cost nearly nothing. However, we are not really interested in such generality. This section specializes to the cases appropriate for discussing quantum operations, namely
$\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ $\;\xleftrightarrow{\;\mathfrak{J}\;}\;$ $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K})),$ (3.1)
${\bm{\theta}}(A)$ $\;\xleftrightarrow{\;\mathfrak{J}\;}\;$ ${\bm{\pi}}(A)\quad(A\colon\mathfrak{B}(\mathcal{H},\mathcal{K})).$ (3.2)
Here, we are using some new notation.
Definition 3.1 ($\bm{\theta}$}).
For $A\colon\mathfrak{B}(\mathcal{H},\mathcal{K})$, ${\bm{\theta}}(A)\in\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{H}))$ is given by
${\bm{\theta}}(A){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}A\,\square\,A^{\dagger}.$ (3.3)
A superoperator of the form ${\bm{\theta}}(A)$ is called a $\bm{\theta}$}-map. $\dashv$
The correspondence ${\bm{\theta}}(A)\;\xleftrightarrow{\;\mathfrak{J}\;}\;{\bm{\pi}}(A)$ is an obvious specialization of $A\overline{B}\;\xleftrightarrow{\;\mathfrak{J}\;}\;A\,\square\,B^{\dagger}$ (2.10). That it suffices to specify $\mathfrak{J}$ in this case follows from the fact that one-dimensional projectors span $\mathfrak{B}(\mathcal{H})$ for any Hilbert space $\mathcal{H}$, in particular $\mathfrak{B}(\mathcal{H},\mathcal{K})$. The point is made more concretely by the polarization identity $A\overline{B}=\tfrac{1}{4}\sum i^{-n}{\bm{\pi}}(A+i^{n}B)$.
The main goal of this section is to show that $\mathfrak{J}$ gives a canonical isometry between the class of CP operators in $\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ and the positive operators in $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$. The fundamental concept of complete positivity is critically introduced in §3.2. To get there, we will be decomposing the spaces in (3.1) layer-by-layer.
3.1 More spaces of operators
We introduce now spaces of operators with special characteristics, along with the attendant notations.
3.1.1 Hermitian operators $\mathrm{SA}(\mathcal{H})$
This is the subclass of $\mathfrak{B}(\mathcal{H})$ defined by
$A\in\mathrm{SA}(\mathcal{H})\;\equiv\;\forall h\colon\mathcal{H},\,\left\langle{h}\middle|{Ah}\right\rangle\in{\mathbb{R}}.$ (3.4)
(‘SA’ stands for ‘self-adjoint’.) $\mathrm{SA}(\mathcal{H})$ is a vector space over $\mathbb{R}$} in its own right, and
$\mathfrak{B}(\mathcal{H})=\mathrm{SA}(\mathcal{H})\oplus_{\mathbb{R}}i\,\mathrm{SA}(\mathcal{H})$ (3.5)
is a direct sum over $\mathbb{R}$} of the hermitian and anti-hermitian subspaces, corresponding to the unique decomposition of any operator into a hermitian and anti-hermitian part.
3.1.2 Positive operators $\mathrm{Pos}(\mathcal{H})$
This is the subclass of $\mathrm{SA}(\mathcal{H})$ defined by
$A\in\mathrm{Pos}(\mathcal{H})\;\equiv\;\forall h\colon\mathcal{H},\;\left\langle{h}\middle|{Ah}\right\rangle\geq 0.$ (3.6)
$\mathrm{Pos}(\mathcal{H})$ is not a vector space, but a closed convex cone. This concept will play a fundamental role in the sequel.
Definition 3.2 (convex set).
A subset $\mathcal{C}$ of a real vector space $\mathcal{V}$ is convex if whenever $x,y$ are in $\mathcal{C}$, the entire line segment $x,y]=\{sx+(1-s)y\,:\,0\leq s\leq 1\}$ is also in $\mathcal{C}$. $\dashv$
Definition 3.3 (cones).
$\mathcal{C}\subseteq\mathcal{V}$ is a cone if whenever $x$ is in $\mathcal{C}$, then so is the entire ray $\{sx\,:\,0<s\}$. $\mathcal{C}$ is a convex cone if it is also convex. An Alternative full characterization is: whenever $x,y$ are in $\mathcal{C}$, then so are all positive linear combinations $\{ax+by\,:\,0<a,b\}$.
A proper cone $\mathcal{C}$ has the additional properties that (i) it is topologically closed, (ii) it has nonempty interior, and (iii) $x,-x\in\mathcal{C}$ implies $x=0$. From (iii) and (i), it easily follows that $\mathcal{V}=\mathcal{C}-\mathcal{C}$, i.e., every $v\colon\mathcal{V}$ is the difference of two vectors in $\mathcal{C}$. $\dashv$
Example 3.1 ($\mathrm{Pos}(\mathcal{K})$ is a proper cone in $\mathrm{SA}(\mathcal{K})$).
(i) $a,b\geq 0$ and $A,B\colon\mathrm{Pos}(\mathcal{K})$ immediately imply $aA+bB\in\mathrm{Pos}(\mathcal{K})$, (ii) $\left\langle{v}\middle|{A_{n}v}\right\rangle\to\left\langle{v}\middle|{Av}\right\rangle$ whenever $A_{n}\to A$, (iii) any sufficiently small perturbation of $\operatorname{\mathbbm{1}}_{\mathcal{K}}$ is positive, (iv) $\left\langle{v}\middle|{Av}\right\rangle\geq 0$ and $\left\langle{v}\middle|{Av}\right\rangle\geq 0$ can hold for all $v$ only if $A=0$. $\dashv$
$\mathrm{SA}(\mathcal{H})$ is the real span, and $\mathfrak{B}(\mathcal{H})$ the complex span, of $\mathrm{Pos}({\mathcal{H}})$. This is so because any hermitian operator is, through its spectral representation, canonically expressed as a difference of two positive operators. We write $A\geq 0$” for “$A$ is positive”. This provides a partial order on $\mathrm{SA}(\mathcal{H})$: $A\leq B$ means $B-A\geq 0$.
3.1.3 Property-preserving superoperators
The adjectives hermitian and positive apply to operators on a Hilbert space, therefore to elements of $\mathfrak{B}(\mathcal{H})$, but generally not to elements of $\mathfrak{B}(\mathcal{H},\mathcal{K})$. Something new, yet related, appears on this level, namely a superoperator may preserve the property of hermiticity or of positivity. We thus define the hermiticity-preserving and monotone superoperators by
For $\Lambda\colon\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$,
$\begin{array}[]{lcl}\Lambda\in\mathsf{HP}(\mathcal{H},\mathcal{K})&\quad\equiv\quad&\quad\Lambda(\mathrm{SA}(\mathcal{H}))\subseteq\mathrm{SA}(\mathcal{K})\\ \Lambda\in\mathsf{Mon}(\mathcal{H},\mathcal{K})&\quad\equiv\quad&\quad\Lambda(\mathrm{Pos}(\mathcal{H}))\subseteq\mathrm{Pos}(\mathcal{K}).\end{array}$ (3.7)
These classes inherit many properties from those of the hermitian and positive operators. For instance, it is easy to see that $\mathsf{HP}(\mathcal{H},\mathcal{K})$ is a real vector space. $\mathsf{Mon}(\mathcal{H},\mathcal{K})$ is a closed convex cone, and since hermitian operators are exactly those which can be expressed as a difference of two positive operators, it is a subset of $\mathsf{HP}(\mathcal{H},\mathcal{K})$. We note that hermiticity-preserving operators can alternatively be defined as those which commute with hermitian conjugation, $\dagger$}.
3.2 $\mathsf{CP}$ is the physically acceptable class of quantum operations
3.2.1 The trouble with $\mathsf{Mon}$
Now we come to the class of superoperators which are at the very center of this investigation. Quantum mechanical density matrices are positive operators, and quantum operations produce density matrix outputs for density matrix inputs. Thus, the superoperator $\Lambda$ representing such an operation with input a state over $\mathcal{H}$ and output a state over $\mathcal{K}$ should be in $\mathrm{Pos}(\mathcal{H},\mathcal{K})$. This is not the only physically-indicated restriction, however.
Example 3.2.
Let $\Lambda$ act on density matrices of a single qubit by transposition in the computational basis; this clearly preserves positivity. Now consider the entangled two-qubit state $\left|{10}\right\rangle+\left|{01}\right\rangle$ (normalization omitted for convenience), with corresponding density matrix
$\rho=\left|{10}\middle\rangle\middle\langle{10}\right|+\left|{01}\middle\rangle\middle\langle{01}\right|+\left|{10}\middle\rangle\middle\langle{01}\right|+\left|{01}\middle\rangle\middle\langle{10}\right|.$ (3.8)
Performing transposition on the first qubit without altering the second delivers a non-positive density matrix:
$(\Lambda\otimes\operatorname{\mathbbm{1}})\rho=\left|{10}\middle\rangle\middle\langle{10}\right|+\left|{01}\middle\rangle\middle\langle{01}\right|+\left|{00}\middle\rangle\middle\langle{11}\right|+\left|{11}\middle\rangle\middle\langle{00}\right|.$ (3.9)
(For instance, take expectation value in the vector state $\left|{00}\right\rangle-\left|{11}\right\rangle$.) $\dashv$
3.2.2 CP makes the world safe for tensor products
Whatever external apparatuses are required to manifest two physically allowable operations $\Lambda$ and $\Gamma$ individually, there is no reason in principle that they should not be able to run simultaneously. This implies that $\Lambda\otimes\Gamma$ should be allowed. (The expression $\Lambda\otimes\Gamma$” implies that $\Lambda$ and $\Gamma$ operate on disjoint sets of degrees of freedom, just as $\mathcal{H}\otimes\mathcal{H}$” implies two distinct copies of $\mathcal{H}$.) Thus, introducing the temporary designation $\mathit{OK}$” for the admissible superoperators, the class of $\mathit{OK}$superoperators must be smaller than that of monotone superoperators. Consider the following postulates for such a class:
  1. O1.
    $\mathit{OK}$ superoperators are monotone;
  2. O2.
    Identity superoperators are $\mathit{OK}$;
  3. O3.
    If $\Lambda$ and $\Gamma$ are $\mathit{OK}$, then $\Lambda\otimes\Gamma$ is $\mathit{OK}$;
  4. O4.
    The class of $\mathit{OK}$ operators is maximal, subject to the preceding.
The issue here is collective; it is not a question of what is an $\mathit{OK}$ map, but, what is an $\mathit{OK}$ class? Identity superoperators $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ are postulated to be $\mathit{OK}$ because they correspond to “nothing happens”. The final postulate merely says that nothing should be disallowed without reason. Fortunately, a unique class is picked out by our postulates, and there is a membership test.
Definition 3.4 ($\mathsf{CP}$).
$\Lambda\colon\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ is completly positive (CP) if
$\forall N\geq 1,\;\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}\in\mathsf{Mon}(\mathcal{H}\otimes{\mathbb{C}}^{N},\mathcal{K}\otimes{\mathbb{C}}^{N})$ (3.10)
The class of CP operators in $\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ is denoted $\mathsf{CP}(\mathcal{H},\mathcal{K})$ (same convention of referring directly to underlying Hilbert spaces as for $\mathsf{HP}$ and $\mathsf{Mon}$). $\dashv$
Proposition 3.1.
The class of $\mathsf{CP}$ maps is the unique class satisfying the $\mathit{OK}$ postulates.
Proof.
O1 : The case $N=1$ in the definition says that $\Lambda$ is monotone.
O2 : obvious.
O3: Assume $\Lambda$ and $\Gamma$ are $\mathsf{CP}$ and compute
$(\Lambda\otimes\Gamma)\otimes\operatorname{\mathbbm{1}}_{\mathcal{O}}=(\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{M}})(\operatorname{\mathbbm{1}}_{\mathcal{H}}\otimes\Gamma)\otimes\operatorname{\mathbbm{1}}_{\mathcal{O}}=(\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{M}}\otimes\operatorname{\mathbbm{1}}_{\mathcal{O}})(\operatorname{\mathbbm{1}}_{\mathcal{H}}\otimes\Gamma\otimes\operatorname{\mathbbm{1}}_{\mathcal{O}}).$ (3.11)
Since permutation of factors in a tensor product is an isometric isomorphism (cf. §2.3.2), $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{M}}\otimes\operatorname{\mathbbm{1}}_{\mathcal{O}}$ and $\operatorname{\mathbbm{1}}_{\mathcal{H}}\otimes\Gamma\otimes\operatorname{\mathbbm{1}}_{\mathcal{O}}$ are monotone by hypothesis, hence so is their composition. Therefore $\Lambda\otimes\Gamma$ is $\mathsf{CP}$.
O4 : Any $\mathit{OK}$ operator must be $\mathsf{CP}$, so $\mathsf{CP}$ is maximal satisfying O1–O3.
By the same token, however, any $\mathit{OK}$ class is a subclass of $\mathsf{CP}$, hence, by maximality, equal to it. ∎
3.3 The central diagram
The spaces we have been discussing are related as $\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})\subset\mathsf{HP}(\mathcal{H},\mathcal{K})$. The point of the Jamiołkowski isomorphism for this is that it will give us isometries according to the following “central” commutative diagram:
$\mathsf{CP}(\mathcal{H},\mathcal{K})$$\mathsf{HP}(\mathcal{H},\mathcal{K})$$\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$$\mathfrak{B}(\mathcal{H},\mathcal{K})$$\mathrm{Pos}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$$\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$$\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$$\scriptstyle{\mathfrak{J}}\scriptstyle{\subset}$$\scriptstyle{\mathfrak{J}}\scriptstyle{\subset}$$\scriptstyle{\mathfrak{J}}$$\scriptstyle{{\bm{\theta}}}$$\scriptstyle{{\bm{\pi}}}\scriptstyle{\subset}\scriptstyle{\subset}$ (3.12)
Note that we write $\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathsf{HP}(\mathcal{H},\mathcal{K})$ even though these are spaces of superoperators $\mathfrak{B}(\mathcal{H})\rightarrow\mathfrak{B}(\mathcal{K})$. All geometrical properties of $\mathsf{CP}(\mathcal{H},\mathcal{K})$ can be studied through $\mathrm{Pos}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$. The rest of this section, and §§4, 5, will show the value of this point of view. The middle items in the rows here are real Hilbert spaces (i.e., over the field $\mathbb{R}$} of scalars). The items on the right are their complexifications. This is familiar for the bottom row through the splitting of an arbitrary operator into hermitian and anti-hermitian parts and then follows for the top row through $\mathfrak{J}$, as soon as the middle $\mathfrak{J}$ arrow is established.
3.4 Terminological dilemmas
The literature traditionally calls monotone operators positive. This usage is adopted from a context ($C^{*}$ algebras) where it is unambiguous, but for us it can cause confusion because $\Lambda\in\mathfrak{B}(\mathfrak{B}(\mathcal{H}))$ might be monotone, and it might be positive (as an operator on the Hilbert space $\mathfrak{B}(\mathcal{H})$). These are very different things. Consistency would recommend “completely monotone” over “completely positive”, but the latter is far too established for that. We escape the dilemma by always using the acronym “CP” rather than the full phrase. Similarly, we will below use “$N$-monotone” instead of the usual “$N$-positive”.
3.5 Demonstration of the central diagram (3.12)
3.5.1 $\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))\;\xleftrightarrow{\;\mathfrak{J}\;}\;\mathsf{HP}(\mathcal{H},\mathcal{K})$
$\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$ is the real span of $\left\{{\bm{\pi}}(A)\,\colon\,A\in\mathfrak{B}(\mathcal{H},\mathcal{K})\right\}$, and if $X\in\mathrm{SA}(\mathcal{H})$, then $AXA^{\dagger}\in\mathrm{SA}(\mathcal{K})$, i.e., ${\bm{\theta}}(A)\in\mathsf{HP}(\mathcal{H},\mathcal{K})$. So, $\mathfrak{J}$ carries $\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$ into $\mathsf{HP}(\mathcal{H},\mathcal{K})$ But, $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$ is the complexification of $\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$, so $\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ is the complexification of $\mathfrak{J}\,[\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))]$. Since no nonreal multiple of a nonzero hermiticity-preserving superoperator can be hermiticity-preserving, $\mathfrak{J}\,[\mathrm{SA}(\mathfrak{B}(\mathcal{H},\mathcal{K}))]=\mathsf{HP}(\mathcal{H},\mathcal{K})$.
3.5.2 $\mathrm{Pos}(\mathfrak{B}(\mathcal{H},\mathcal{K}))\;\xleftrightarrow{\;\mathfrak{J}\;}\;\mathsf{CP}(\mathcal{H},\mathcal{K})$
Establishing the positive-CP isomorphism will be more work and involves some intermediate notions.
Definition 3.5 (NN-monotone, rank-NN positive).
$\Lambda\in\left({\mathfrak{B}(\mathcal{K})}\leftarrow{\mathfrak{B}(\mathcal{H})}\right)$ is $N$-monotone if $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}$ is monotone. So, 11-monotone is the same as simply monotone.
On the other side, $\Lambda\in\mathfrak{B}(\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right))$ is rank-$N$-positive if $\left\langle{T}\middle|{\Lambda T}\right\rangle\geq 0$ for all $T$ of rank not exceeding $N$. $\dashv$
Whatever intrinsic interest these may hold, for us they are purely instrumental, so the reader should feel free to forget these terms after this section. With them, we have an expansion of the information in the central diagram, presented now in tabular form:
${\mathfrak{B}(\mathcal{H})}\rightarrow{\mathfrak{B}(\mathcal{K})}$ $\mathfrak{B}({\mathcal{H}}\rightarrow{\mathcal{K}})$
hermiticity-preserving hermitian
monotone rank-11-positive
NN-monotone rank-NN-positive
CP positive
Since CP is equivalent to ($\forall N$, NN-monotone) and positive is equivalent to ($\forall N$, rank-NN-positive), “CP $\xleftrightarrow{\mathfrak{J}}$ positive” follows immediately from “$N$-monotone $\xleftrightarrow{\mathfrak{J}}$ rank-NN-positive”.
We start with the special case “monotone $\xleftrightarrow{\mathfrak{J}}$ rank-11-positive”, using the correspondence
${\bm{\pi}}(k)\,\square\,{\bm{\pi}}(h)={\bm{\pi}}(k\overline{h})\;\xleftrightarrow{\;\mathfrak{J}\;}\;{\bm{\theta}}(k\overline{h})={\bm{\pi}}(k)\overline{{\bm{\pi}}(h)}.$ (3.13)
The first equality here is from ${\bm{\pi}}(k)A{\bm{\pi}}(h)=k\left\langle{k}\middle|{Ah}\right\rangle\overline{h}=k\overline{h}\left\langle{k\overline{h}}\middle|{A}\right\rangle$ [cf. (2.16)]. Then, we have
$\begin{split}\left\langle{k}\middle|{(\Lambda\cdot{\bm{\pi}}(h))k}\right\rangle&=\left\langle{{\bm{\pi}}(k)}\middle|{\Lambda\cdot{\bm{\pi}}(h)}\right\rangle\\ &=\langle{{\bm{\pi}}(k)\overline{{\bm{\pi}}(h)}}\mid{\Lambda}\rangle\\ &=\left\langle{{\bm{\pi}}(k\overline{h})}\middle|{\Lambda}\right\rangle\\ &=\left\langle{k\overline{h}}\middle|{\Lambda\cdot k\overline{h}}\right\rangle\end{split}$ (3.14)
$\Lambda$ is monotone if and only if the LHS is non-negative for all $h$ and $k$. On the other hand, $k\overline{h}$ ranges over the rank-1 operators in $\left({\mathcal{K}}\leftarrow{\mathcal{H}}\right)$ as $h\colon\mathcal{H}$ and $k\colon\mathcal{K}$ are varied.
Now for the final step.
Lemma 3.2.
For $\Lambda\in\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$, the following conditions are equivalent.
  1. 1.
    $\Lambda$ is rank-$N$-positive
  2. 2.
    $\Lambda\otimes{\bm{\pi}}[\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}]$ is rank-1-positive
  3. 3.
    $\mathfrak{J}\Lambda$ is $N$-monotone
Proof.
(1) $\Leftrightarrow$ (2): Let $\{e_{i}\}$ be an ONB of ${\mathbb{C}}^{N}$, and define for $h_{i}\colon\mathcal{H}$ and $k_{i}\colon\mathcal{K}$,
$\Gamma{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sum_{i=1}^{N}k_{i}\overline{h_{i}},\quad\Phi{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sum_{i=1}^{N}k_{i}\otimes e_{i},\quad\Psi{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sum_{i=1}^{N}h_{i}\otimes e_{i}.$ (3.15)
As the $h_{i}$ and $k_{i}$ range over their domains, $\Gamma$ ranges over the rank-$N$ operators $\mathcal{H}\rightarrow\mathcal{K}$} and $\Phi\overline{\Psi}$ over the rank-$1$ operators $\mathcal{H}\otimes{\mathbb{C}}^{N}\rightarrow\mathcal{K}\otimes{\mathbb{C}}^{N}$. Now,
$\begin{split}\langle{\Phi\overline{\Psi}}\mid{\Lambda\otimes{\bm{\pi}}[\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}]\cdot\Phi\overline{\Psi}}\rangle&=\sum_{i,j,l,m}\left\langle{k_{i}\overline{h_{j}}}\middle|{\Lambda\cdot k_{l}\overline{h_{m}}}\right\rangle\left\langle{e_{i}\overline{e_{j}}}\middle|{\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}}\right\rangle\left\langle{\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}}\middle|{e_{l}\overline{e_{m}}}\right\rangle\\ &=\sum_{i,l}\left\langle{k_{i}\overline{h_{i}}}\middle|{\Lambda\cdot k_{l}\overline{h_{l}}}\right\rangle\\ &=\left\langle{\Gamma}\middle|{\Lambda\Gamma}\right\rangle.\end{split}$ (3.16)
(2) $\Leftrightarrow$ (3): Remember that (2.11) $\mathfrak{J}$ distributes over $\otimes$, so
$\mathfrak{J}(\Lambda\otimes{\bm{\pi}}[\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}])=\mathfrak{J}\Lambda\otimes\mathfrak{J}{\bm{\pi}}[\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}]=\mathfrak{J}\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}.$ (3.17)
Now, apply “monotone $\xleftrightarrow{\mathfrak{J}}$ rank-11-positive”. ∎
4 Properties of CP maps via Jamiołkowsi
Now that we know that the Jamiołkowsi isomorphism is an isometry as indicated in the central diagram (3.12), we can proceed to deduce geometric properties of $\mathsf{CP}(\mathcal{H},\mathcal{K})$ from those of $\mathrm{Pos}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$. This is what we shall do in this section.
4.1 Kraus decomposition is extremal decomposition
Example 3.1 shows that $\mathrm{Pos}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$ is a proper cone in $\mathrm{SA}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{N}))$, so the same is true of $\mathsf{CP}(\mathcal{H},\mathcal{N})$ as a subset of $\mathsf{HP}({\mathcal{H}},{\mathcal{N}})$. A map in $\mathsf{CP}(\mathcal{H},\mathcal{N})$ may be decomposable as a sum of other CP maps. The questions addressed here is whether there are indecomposable CP maps and whether any CP map can be decomposed into a sum of indecomposables. Rather than “indecomposable”, a more geometric terminology will be preferred. The main contention of this subsection was noted by Lindblad (Lemma 4 of Ref. [2], relying on operator-algebraic representation theory).
Definition 4.1 (extreme point, extreme vector, extreme CP map, extremal decomposition).
A point of a convex set $\mathcal{C}$ is extreme if it cannot be expressed as a convex combination $\lambda y+(1-\lambda)z$, $0<\lambda<1$ for distinct points $y,z$ in $\mathcal{C}$.
For a convex cone $\mathcal{C}$, a slight variation on this concept is more useful, because the only extreme point (if any) is the origin. Vector $x$ in $C$ is an extreme vector, and the ray it is on is an extreme ray, if it cannot be expressed as a sum $y+z$ of linearly independent nonzero $y,z$ in $\mathcal{C}$.
Extremal decomposition refers to either: an expression of a vector in a convex cone as a sum of extreme vectors, or an expression of a point in a convex set as a convex combination of extreme points. Context should disambiguate.
Since the term “vector” can be confusing in our context, we will refer to elements of an extreme ray of $\mathsf{CP}(\mathcal{H},\mathcal{K})$ simply as extreme CP maps. $\dashv$
For any Hilbert space $\mathcal{M}$, every operator in $\mathrm{Pos}(\mathcal{M})$ is decomposable into a sum of extreme vectors, and the latter are the rank-1 elements ${\bm{\pi}}(m)$. We prove this momentarily, but first observe that — via $\mathfrak{J}$ — it implies (i) the extreme vectors in the cone $\mathsf{CP}(\mathcal{H},\mathcal{K})$ are precisely those of form ${\bm{\theta}}(A)$, (ii) any $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ can be expressed as $\sum{\bm{\theta}}(A_{i})$, for some collection of operators $A_{i}\colon\mathfrak{B}(\mathcal{H},\mathcal{K})$, (iii) the $A_{i}$ may even be chosen mutually orthogonal. Such an extremal decomposition of a CP superoperators is known as a Kraus decomposition.
That every operator in $\mathrm{Pos}(\mathcal{M})$ can be written as a sum $\sum{\bm{\pi}}(m_{i})$ is a simple matter of finding eigenvectors. That ${\bm{\pi}}(m)$ is actually extreme follows immediately from
Lemma 4.1.
If $S,T,W\colon\mathrm{Pos}(\mathcal{H})$ satisfy $S+T=W$, then $\operatorname{ran}S+\operatorname{ran}T=\operatorname{ran}W$.
Proof.
We rely on the basic fact that for any positive operator $M$,
$\ker A=\left\{v\colon\mathcal{H}\,\colon\,\left\langle{v}\middle|{Av}\right\rangle=0\right\}.$ (4.1)
$\ker W=\ker S\cap\ker T$ follows immediately. Now use $(\operatorname{ran}A)^{\perp}=\ker A^{\dagger}$, valid for any operator, and $(\operatorname{ran}S)^{\perp}\cap(\operatorname{ran}T)^{\perp}=(\operatorname{ran}S+\operatorname{ran}T)^{\perp}$. ∎
Our proof of the existence of extremal decomposition relied on the existence of a complete set of eigenvectors. This is certainly a good argument, but one more in keeping with the geometric flavor can be made by appeal to the (finite-dimensional) Krein-Milman theorem which says that extremal decompositions always exist for a compact (i.e., closed and bounded) convex set. Every ray in the cone $\mathrm{Pos}(\mathcal{H})$ intersects the affine hyperplane of trace-1 operators, and that intersection is compact. Thus, for a normalized multiple of any positive operator a decomposition into a convex combination of extreme trace-1 positive operators exists by Krein-Milman, and this can be used to obtain a decomposition of the original operator.
4.2 $\bm{\theta}$ is a functor from the world of closed quantum systems to that of open quantum systems
In this section, we study the collection of all CP maps over all pairs of underlying Hilbert spaces — the category $\mathsf{CP}$. The purpose is to open a broader perspective and demonstrate the utility of our Jamiołkowski approach in a context which is not geometrical, where composition is paramount.
Consider the collection of all finite-dimensional Hilbert spaces (the objects of the category) and linear maps between them (the morphisms). We call this category $\mathsf{Hilb}$. To each ordered pair $(\mathcal{H},\mathcal{K})$ of Hilbert spaces there corresponds the collection $\mathfrak{B}(\mathcal{H},\mathcal{K})$ of linear maps from $\mathcal{H}$ to $\mathcal{K}$. If $A\colon\mathfrak{B}(\mathcal{H},\mathcal{K})$ and $B\colon\mathfrak{B}(\mathcal{K},\mathcal{N})$, they can be composed as $BA$ to yield a map in $\mathfrak{B}(\mathcal{H},\mathcal{N})$. Further, for each $\mathcal{H}$, there is an identity $\operatorname{\mathbbm{1}}_{\mathcal{H}}$ in $\mathfrak{B}(\mathcal{H},\mathcal{H})$ such that $A\operatorname{\mathbbm{1}}_{\mathcal{H}}=A$ or $\operatorname{\mathbbm{1}}_{\mathcal{H}}B=B$ whenever these compositions are well-typed. These simple structural properties license $\mathsf{Hilb}$ as a category. Note that the objects do not really “do” anything except identify which pairs of morphisms need to be composable.
More interesting are that CP maps also form a category, $\mathsf{CP}$, and that $\bm{\theta}$} is a functor from $\mathsf{Hilb}$ to $\mathsf{CP}$, that is, a map which respects the categorical structure. Let us demonstrate these. Suppose $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\Gamma\colon\mathsf{CP}(\mathcal{K},\mathcal{N})$ given. Then $\Gamma\Lambda$ is a linear map $\mathfrak{B}(\mathcal{H},\mathcal{N})$ and we aim to show that it is CP following (3.10). Now, $\Lambda$ and $\Gamma$ are monotone — they preserve the property of being positive. It is then immediate that their composite is also monotone. So, if $\Lambda$ and $\Gamma$ are $N$-montone, so is the LHS of
$(\Gamma\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})})(\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})})=\Gamma\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}.$ (4.2)
Monotonicity of the RHS for all $N$ now says that $\Gamma\Lambda\in\mathsf{CP}(\mathcal{H},\mathcal{N})$. That $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H},\mathcal{K})}$ is an identity morphism in $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is obvious, so $\mathsf{CP}$ is a category, as claimed. As noted earlier, objects of a category are not required to do anything besides serve as labels. $\mathsf{CP}(\mathcal{H},\mathcal{K})$ consists of linear maps $\mathfrak{B}(\mathfrak{B}(\mathcal{H}),\mathfrak{B}(\mathcal{K}))$ so an immediate impulse is to consider $\mathfrak{B}(\mathcal{H})$ and $\mathfrak{B}(\mathcal{K})$ to be the source and target objects. We can as well consider the underlying Hilbert spaces (here, $\mathcal{H}$ and $\mathcal{K}$) to be the objects. The latter view is implicit in the notation $\mathsf{CP}(\mathcal{H},\mathcal{K})$, and is the view taken in this subsection.
Next, we turn to $\bm{\theta}$} (3.3). It maps $\mathfrak{B}(\mathcal{H},\mathcal{K})$ into $\mathsf{CP}(\mathcal{H},\mathcal{K})$ in such a way that whenever $BA$ is a valid composition, ${\bm{\theta}}(BA)={\bm{\theta}}(B){\bm{\theta}}(A)$, and ${\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$. These are simple to verify and are exactly what is required to show that $\bm{\theta}$} is a functor $\mathsf{Hilb}\rightarrow\mathsf{CP}$. Actually, that is a bit of a distortion. A functor consists of both a map on objects and a map on morphisms. Here, the object map is implicitly the identity. $\bm{\theta}$} is very nearly an embedding of $\mathsf{Hilb}$ into $\mathsf{CP}$. Its failure to be such consists of the fact that it conflates maps differing only by a phase factor.
Both $\mathsf{Hilb}$ and $\mathsf{CP}$ are closed under tensor product. For $\mathsf{CP}$, this property was shown in §3.2. The terminology is that they are monoidal categories under $\otimes$. Furthermore, $\bm{\theta}$} is a monoidal functor by virtue of the fact that it respects tensor products:
${\bm{\theta}}(A\otimes B)={\bm{\theta}}(A)\otimes{\bm{\theta}}(B).$ (4.3)
This is immediately verified.
The above discussion showed in passing that $\mathsf{Mon}$, defined in an obvious manner, is also a category. We can rephrase what was shown in §3.2 as: $\mathsf{CP}$ is the largest subcategory of $\mathsf{Mon}$ compatible with tensor product.
4.3 Further examples
4.3.1 When are Kraus sums equal?
What are necessary and sufficient conditions on $A_{i}$, $B_{j}$ such that $\sum_{i}{\bm{\theta}}(A_{i})=\sum_{j}{\bm{\theta}}(B_{j})$? $\mathfrak{J}$ allows us to translate this to the completely equivalent question, when is $\sum_{i}{\bm{\pi}}(h_{i})$ equal to $\sum_{j}{\bm{\pi}}(k_{j})$?
Lemma 4.2.
Suppose $\left\{x_{i}\,\colon\,i=1,\ldots,n\right\}$ are linearly independent vectors in $\mathcal{H}$. Then $\sum_{\alpha=1}^{m}{\bm{\pi}}(y_{\alpha})=\sum_{i=1}^{n}{\bm{\pi}}(x_{i})$ if and only if $y_{\alpha}=\sum_{i}U_{\alpha}^{i}x_{i}$ for unique numerical vectors $U^{i}$ which are orthonormal in the sense that
$(U^{j})^{\dagger}\cdot U^{k}=\sum_{\alpha}\overline{U_{\alpha}^{j}}{U_{\alpha}^{k}}=\delta^{jk}.$ (4.4)
When this holds, the matrix $\begin{pmatrix}U^{1}\cdots U^{m}\end{pmatrix}$ can be completed to a unitary matrix $\tilde{U}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\begin{pmatrix}U^{1}\cdots U^{n}\end{pmatrix}$ such that $y=\tilde{U}\cdot\tilde{x}$, where $\tilde{x}=(x_{1},\cdots,x_{m},0,\cdots,0)^{T}$ is $x$ padded out with nmn-m zero vectors.
Proof.
If the hypothesis holds, it is necessary that each $k_{\alpha}$ be in the span of $\{x_{1},\ldots,x_{n}\}$ (cf. Lemma 4.1). Hence $y=U\cdot x$ for some matrix of coefficients $U$, and $\sum_{j}x_{j}\overline{x_{j}}=\sum_{\alpha}{\bm{\pi}}(y_{\alpha})=\sum_{jk}\left(U^{j}\cdot\overline{U^{k}}\right)x_{j}\overline{x_{k}}$. Since the $x_{i}$ are linearly independent, (4.4) follows. Sufficiency is clear.
Now, given $y=Ux$, it follows that $C^{\dagger}y=C^{\dagger}Ux=\sum_{i}(C^{\dagger}U^{i})x_{i}$. So, if $U^{m+1},\ldots,U^{n}$ are chosen to fill out an orthonormal basis $U^{1},\ldots,U^{n}$ of ${\mathbb{C}}^{n}$, then
$(U^{i})^{\dagger}y=\begin{cases}x_{i}&1\leq i\leq m\\ 0&m<i\leq n,\end{cases}$ (4.5)
which is to say, $\tilde{U}^{\dagger}y=\tilde{x}$. ∎
Proposition 4.3.
$\sum_{i=1}^{n}{\bm{\pi}}(h_{i})=\sum_{i=1}^{n+m}{\bm{\pi}}(k_{j})$ if and only if $\tilde{h}=U{k}$ for some unitary matrix $U$, where $\tilde{h}$ is $h$ padded out with $m$ zero entries.
Proof.
$\sum{\bm{\pi}}{h_{i}}$ is a positive operator. Let $\sum{\bm{\pi}}{m_{j}}$ be a decomposition with mutually orthogonal $m_{j}$ and apply the Lemma twice, to go from the $k$’s to the $m$’s, then to the $h$’s. ∎
Remark 4.1.
The above is essentially the same as Theorem 2.6 of Ref. [15] (p. 103), with a little more attention to the tricky part about padding with zeros. Theorem 8.2 loc. cit. (p. 372) uses the earlier theorem to prove when Kraus sums are equal. Our main point is simply that the latter result follows immediately from Lemma 4.2. The result will actually be used in §10.4.
4.3.2 Partial trace is CP
First we make an auxiliary observation: using eigenvectors, a tensor product of positive operators can be written as a sum of operators of the form ${\bm{\pi}}[v]\otimes{\bm{\pi}}[w]$, but
${\bm{\pi}}[v]\otimes{\bm{\pi}}[w]={\bm{\pi}}[v\otimes w].$ (4.6)
Therefore, a tensor product of positive operators is positive. Via Jamiołkowski, in particular (2.11) we conclude that a tensor product of CP maps is CP. Therefore, if $\operatorname{\mathrm{Tr}}_{\mathcal{K}}\colon\mathfrak{B}(\mathcal{K})\rightarrow{\mathbb{C}}$ is CP, then so is the partial trace
$\operatorname{\mathrm{Tr}}_{\mathcal{K}}\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}\colon\mathfrak{B}(\mathcal{K})\otimes\mathfrak{B}(\mathcal{H})\rightarrow\mathfrak{B}(\mathcal{H})(={\mathbb{C}}\otimes\mathfrak{B}(\mathcal{H})).$ (4.7)
But,
$\operatorname{\mathrm{Tr}}_{\mathcal{K}}=\overline{\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K})}}=\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}})}\overline{\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K})}}\;\xmapsto{\mathfrak{J}}\;\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}})}\,\square\,{\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K})}}=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K},{\mathbb{C}})}=\operatorname{\mathbbm{1}}_{\mathcal{K}^{*}}.$ (4.8)
The last expression being positive shows that $\operatorname{\mathrm{Tr}}_{\mathcal{K}}$ is CP.
4.3.3 $\mathsf{CP}({\mathbb{C}},\mathcal{H})\cong\mathrm{Pos}(\mathcal{H})\cong\mathsf{CP}(\mathcal{H},{\mathbb{C}})$
The isomorpisms in the title follow directly from the relevant special cases of the Jamiołkowsi isomorphism: $\mathsf{CP}({\mathbb{C}},\mathcal{H})\xleftrightarrow{\mathfrak{J}}\mathrm{Pos}(\mathfrak{B}({\mathbb{C}},\mathcal{H}))=\mathrm{Pos}(\mathcal{H})$ and $\mathsf{CP}(\mathcal{H},{\mathbb{C}})\xleftrightarrow{\mathfrak{J}}\mathrm{Pos}(\mathfrak{B}(\mathcal{H},{\mathbb{C}}))=\mathrm{Pos}(\mathcal{H}^{*})\cong\mathrm{Pos}(\mathcal{H})$. In writing these, we are using the identifiation of $h\colon\mathcal{H}$ with the map $c\mapsto ch\colon{\mathbb{C}}\rightarrow\mathcal{H}$, which gives $\mathfrak{B}({\mathbb{C}},\mathcal{H})\cong\mathcal{H}$. Also, $\mathcal{H}\otimes\mathcal{H}^{*}\cong\mathcal{H}\otimes\mathcal{H}^{*}$.
But what, concretely, is the CP map corresponding to $A\colon\mathrm{Pos}(\mathcal{H})$ here? To make sense of this, we need to recognize that, regarding $h\colon\mathcal{H}$ as an element of $\mathfrak{B}({\mathbb{C}},\mathcal{H})$, $v^{\dagger}=\overline{v}$ and elements of $\mathfrak{B}({\mathbb{C}})$ are just (multiplication by) complex numbers, hence can be floated out of any expression. Then,
$\begin{split}{\bm{\theta}}(h)&=h\,\square\,h^{\dagger}=h\,\square\,\overline{h}=h\overline{h}={\bm{\pi}}(h),\\ {\bm{\theta}}(\overline{h})&=\overline{h}\,\square\,\overline{h}^{\dagger}=\overline{h}\,\square\,{h}=\overline{{\bm{\pi}}(h)}\end{split}$ (4.9)
Thus, the elements of $\mathsf{CP}({\mathbb{C}},\mathcal{H})$ and $\mathsf{CP}(\mathcal{H},{\mathbb{C}})$ corresponding to $A\colon\mathrm{Pos}(\mathcal{H})$ are, respectively, $A$ itself and the linear map $B\mapsto\left\langle{A}\middle|{B}\right\rangle$. In particular (take $A=\operatorname{\mathbbm{1}}_{\mathcal{H}}$), this again shows that trace is CP; compare example 4.3.2. The first displayed equation also shows that the formula ${\bm{\theta}}(A){\bm{\pi}}(h)={\bm{\pi}}(Ah)$, is actually an instance of functoriality.
These results say something interesting about the “world” (i.e., category) of CP maps. Namely, density operators, the things on which CP maps operate, thereby acquiring meaning, are themselves CP maps. Thus, the action of a CP map is expressible purely as a composition in $\mathsf{CP}$, e.g., $\Lambda\circ\rho$. This — via the functor $\bm{\theta}$} — mirrors the corresponding situation in the world $\mathsf{Hilb}$ of closed quantum systems, where we have $\mathfrak{B}({\mathbb{C}},\mathcal{H})\cong\mathcal{H}$.
4.3.4 Isomorphisms in CP ==\, $\bm{\theta}$(isomorphisms in $\mathsf{Hilb}$)
cf. Thm. 3.4.1 of [4] An isomorphism is a morphism with both a left and right inverse (in which case those two are the same). It is easy to see that if $A:\mathfrak{B}(\mathcal{H},\mathcal{K})$ is invertible, then ${\bm{\theta}}(A)^{-1}={\bm{\theta}}(A^{-1})$. (This is a general property of functors, in fact.)
For the converse, suppose that $\Lambda=\sum_{i}{\bm{\theta}}(A_{i})$, $A_{i}\colon\mathfrak{B}(\mathcal{H},\mathcal{K})$ has a two-sided inverse $\Gamma=\sum_{j}{\bm{\theta}}(B_{j})$ with $B_{j}\colon\mathfrak{B}(\mathcal{K},\mathcal{H})$. That is, $\Gamma\Lambda=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ and $\Lambda\Gamma=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K})}$. Then,
$\sum_{ij}{\bm{\theta}}(B_{j}A_{i})=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})},\qquad\sum_{ij}{\bm{\theta}}(A_{i}B_{j})=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K})}.$ (4.10)
But the identities $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}={\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})$ and $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ are extreme maps in $\mathsf{CP}(\mathcal{H})$ and $\mathsf{CP}(\mathcal{K})$, respectively. Therefore, $A_{i}B_{j}=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{K})}$ and $B_{j}A_{i}=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ for all $i,j$. Equivalently, $B_{j}=A_{i}^{-1}$. This implies that every $A_{i}\propto A$ for some $A$, and $B_{j}\propto A^{-1}$, i.e., $\Lambda={\bm{\theta}}(A)$, $\Gamma={\bm{\theta}}(A^{-1})$.
4.4 Trace-related properties
If a CP map is to preserve the normalization of density operators, it must be trace-preserving in the sense of
Definition 4.2 (trace-preserving, trace-non-increasing).
$\Lambda\colon\mathsf{HP}(\mathcal{H},\mathcal{K})$ is trace-non-increasing, respectively trace-preserving when $\operatorname{\mathrm{Tr}}\Lambda\rho\leq\operatorname{\mathrm{Tr}}\rho$, respectively $\operatorname{\mathrm{Tr}}\Lambda\rho=\operatorname{\mathrm{Tr}}\rho$, for every $\rho\colon\mathrm{Pos}(\mathcal{H})$. Many authors call trace-non-increasing CP maps quantum operations and trace-preserving ones quantum channels. $\dashv$
The weaker property of being trace-non-increasing is definitely of interest in quantum physics, see, e.g. Refs. [12, 9, 15].
Since if $A\colon\mathfrak{B}(\mathcal{H})$, $\operatorname{\mathrm{Tr}}A=\left\langle{\operatorname{\mathbbm{1}}_{\mathcal{H}}}\middle|{A}\right\rangle$, a superoperator $\Lambda\colon\mathsf{HP}(\mathcal{H},\mathcal{K})$ is trace-non-increasing precisely when
$\Lambda^{\dagger}\operatorname{\mathbbm{1}}_{\mathcal{K}}\leq\operatorname{\mathbbm{1}}_{\mathcal{H}},$ (4.11)
and trace-preserving when equality holds there. Small enough multiples of any CP (or even monotone) $\Lambda$ are trace-non-increasing. In fact, $\|\Lambda\|\leq 1$ (operator norm) suffices. On the other hand, there may be no multiple which is trace-preserving. For instance, if $\Lambda={\bm{\theta}}(A)$, there will be one only if $A$ is proportional to a unitary.
5 Markovian CP evolution and tangent geometry of $\mathsf{CP}(\mathcal{H})$
This Section turns to the GKSL problem with the tools we have developed. First we bring to mind salient features of continuous unitary evolution, appropriate for closed systems. This will serve as a reference/contrast/inspiration for the discussion of CP evolution to follow. There is an issue of Markovianity which arises for non-stationary evolution (i.e., the evolution itself is not time-translation invariant). The tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ to $\mathsf{CP}(\mathcal{H})$ is brought into the picture and it is shown that a continuously varying generator in a subclass thereof (the $L$-cone) gives Markovian CP evolution. The proof, with minor qualification, works in infinite dimensions. Finally, we show that the $L$-cone is all of $\mathsf{cp}^{+}(\mathcal{H})$. This part relies on $\mathfrak{J}$, hence is inherently finite-dimensional.
Here is a capsule summary, using some terminology that will be explained in the body.
  1. 1.
    In order for $\Lambda(t)$ to generate a right-differentiable Markovian CP evolution (generalizing CP semigroups),
    • it is necessary that $\Lambda(t)$ be in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$,
    • it is sufficient that $\Lambda(t)$ also be piecewise continuous.
  2. 2.
    The tangent space $\mathsf{cp}(\mathcal{H})$ is that part of the tangent cone consisting of operators tangent to $\mathsf{CP}(\mathcal{H})$ both “forward and backward”. These maps correspond to reversible evolutions.
  3. 3.
    $\mathsf{cp}^{+}(\mathcal{H})$ consists of the operators in $\mathsf{HP}(\mathcal{H})$ which can be parametrized (nonuniquely) by $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ as $L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$.
  4. 4.
    Linear operators (Lindblad parametrizers) exist which (i) extract $L$-parametrizations, (ii) map $\mathsf{cp}^{+}(\mathcal{H})$ into $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$, and (iii) map the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ into $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. These provide a parametrization varying continuously with the generator.
  5. 5.
    Conventional parametrization results from Kraus decomposition of $\Psi$ and imposition of trace-preservation.
5.1 Comparison case: $\mathsf{U}(\mathcal{H})$ and unitary evolution
$\mathsf{U}(\mathcal{H})$ here denotes the group of unitary operators on $\mathcal{H}$. Consider a family $\left\{U(t,s)\colon\mathsf{U}(\mathcal{H})\,\colon\,t\geq s\right\}$ for $s,t$ in some interval including zero satisfying the conditions
$t_{2}>t_{1}>t_{0}\;\Rightarrow\;U(t_{2},t_{0})=U(t_{2},t_{1})U(t_{1},t_{0}),$ (5.1a)
$U(t,s)\text{ is continuous in both arguments}.$ (5.1b)
(5.1a) is a kind of Markovianity condition, insofar as it says that to evolve $\psi(t_{0})$, the state at time $t_{0}$, to $t_{2}$, it suffices to know the state at any intermediate time $t_{1}$; the history whereby $\psi(t_{1})$ is reached is irrelevant. Thus, we call such a family a Markovian (unitary) evolution. Note also that the conditions imply that $U(t,t)=\operatorname{\mathbbm{1}}_{\mathcal{H}}$.
We can reduce this family to another indexed by a single time. Let $V(t)$ be any continuous curve in $\mathsf{U}(\mathcal{H})$ satisfying $V(0)=\operatorname{\mathbbm{1}}_{\mathcal{H}}$. Now, defining $U(t,0){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}V(t)$, we recover our Markovian evolution
$U(t,s)=U(t,0)U(s,0)^{-1}=V(t)V(s)^{-1}.$ (5.2)
The defining conditions are immediately verified. For this to work, it is crucially important that $\mathsf{U}(\mathcal{H})$ is a group so that first $V(s)^{-1}$ and then $V(t)V(s)^{-1}$ are unitary.
In keeping with the familiar idea that the derivative of a time evolution operator corresponds to a Hamiltonian, it makes sense to demand that $U(t,s)$ be not merely continuous, but everywhere differentiable, or even continuously differentiable, at least in $t$. Thus, differentiating $U(t,0)=U(t,s)U(s,0)$ with respect to $t$ at $t=s$ gives $\tfrac{d}{dt}V(t)=-iH(t)V(t)$, where
$iH(s)=\lim_{\epsilon\downarrow 0}\tfrac{1}{\epsilon}\left(U(s+\epsilon,s)-\operatorname{\mathbbm{1}}_{\mathcal{H}}\right).$ (5.3)
We have been explicit here that this is a right derivative, because we have defined $U(t,s)$ only for $t\geq s$. The formula (5.3) rearranges to $U(s+\epsilon,s)=\operatorname{\mathbbm{1}}_{\mathcal{H}}+\epsilon(-iH(s))+o(\epsilon)$ for $\epsilon>0$. In geometrical terminology, we might phrase this as: iH(s)-iH(s) is forward tangent to $\mathsf{U}(\mathcal{H})$.
However, as we have already emphasized, $\mathsf{U}(\mathcal{H})$ is a group, so $U(s+\epsilon,s)^{-1}=\operatorname{\mathbbm{1}}_{\mathcal{H}}-\epsilon(-iH(s))+o(\epsilon)$ is also in $\mathsf{U}(\mathcal{H})$. In other words, iH(s)-iH(s) is simply tangent to $\mathsf{U}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}}$, in the ordinary sense; $\operatorname{\mathbbm{1}}_{\mathcal{H}}+\epsilon(-iH(s))$ is within $o(\epsilon)$ regardless of the sign of $\epsilon$. To put it plainly, $iH(s)$ lies in the tangent space $\mathfrak{u}(\mathcal{H})=i\cdot\mathrm{SA}(\mathcal{H})$. This may seem much ado about nothing; the point is that $\mathsf{CP}(\mathcal{H})$ is very different from $\mathsf{U}(\mathcal{H})$ in this respect, as will soon be made clear.
Here are two more related respects in which differentiable Markovian unitary evolution differs from its CP counterpart. Any $U^{\prime}\in\mathsf{U}(\mathcal{H})$ can be reached by such an evolution. All that is required is a differentiable curve $V(t)$ connecting the identity to $U^{\prime}$. This property, too, fails for CP evolutions and is also connected to the fact that $\mathsf{CP}(\mathcal{H})$ is not a group. And, the reversed evolution obtained by changing $H(t)$ to H(Tt)-H(T-t) is also Markovian unitary.
5.2 Half-curves in $\mathsf{CP}(\mathcal{H})$
Now we turn to the object of real interest, $\mathsf{CP}(\mathcal{H})$. Whereas $\mathsf{U}(\mathcal{H})$ is a group and a smooth submanifold of $\mathfrak{B}(\mathcal{H})$, $\mathsf{CP}(\mathcal{H})$ is only a monoid, and a closed convex cone. [According to algebraists and computer scientists, a semigroup is a structure with an associative binary operation and a monoid is a semigroup with an identity element. There is evident tension with the tradition of “quantum dynamical semigroups”.]
5.2.1 Markovian CP evolution
Again, we consider a family $\mathcal{E}(t,s)\colon\mathsf{CP}(\mathcal{H})$ for $0\leq t,s<\omega$ (possibly $\omega=\infty$), satisfying the conditions (5.1a,5.1b). Since there are noninvertible idempotent CP maps, this is not enough to guarantee $\mathcal{E}(t,t)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$. So, we must require also $\mathcal{E}(0,0)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$. With this addition, $\mathcal{E}(t,0)$ is invertible in $\mathfrak{B}(\mathfrak{B}(\mathcal{H}))$ (not in $\mathsf{CP}(\mathcal{H})$!) for all $t\geq 0$, and therefore $\mathcal{E}(t,t)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$.
Proof.
Suppose $t_{0}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\inf\{t\,:\,\mathcal{E}(t,0)\text{ is not invertible}\}$ exists. By continuity in the first and second arguments, respectively, both $\mathcal{E}(t_{0}+\epsilon,t_{0})$ and $\mathcal{E}(t_{0},t_{0}-\epsilon)$ are invertible for some $\epsilon>0$. But, this implies that $\mathcal{E}(t_{0}+\epsilon,0)=$ $\mathcal{E}(t_{0}+\epsilon,t_{0})\mathcal{E}(t_{0},t_{0}-\epsilon)\mathcal{E}(t_{0}-\epsilon,0)$ is also invertible. ∎
A consequence is that, in contrast to the case for unitary evolutions, there are many CP maps which cannot be reached by such a continuous Markovian evolution, because not all CP maps are invertible.
We now we add the assumption that $\mathcal{E}(t,s)$ is continuously differentiable in $t$, and find, just as before,
$\tfrac{\partial}{\partial t}\mathcal{E}(t,0)=\Lambda(t)\mathcal{E}(t,0).$ (5.4)
Invertibility of $\mathcal{E}(t,0)$ secures both existence and uniqueness of $\Lambda(t)$, which, by assumption, is continuous. Multiplying on the right by $\mathcal{E}(s,0)^{-1}$ (for $t\geq s$) yields
$\tfrac{\partial}{\partial t}\mathcal{E}(t,s)=\Lambda(t)\mathcal{E}(t,s).$ (5.5)
Considering this for $t\downarrow s$, we see that $\Lambda$ is forward-tangent to $\mathsf{CP}(\mathcal{H})$ at the identity, a concept which we now formally define.
5.2.2 Tangency and forward tangency
Definition 5.1 (forward-tangent, tangent cone).
If $x$ is a point in a subset $\mathscr{S}$ of a normed space $\mathscr{X}$, the vector $v\colon\mathscr{X}$ is forward-tangent to $\mathscr{S}$ at $x$ precisely when
$x+tv+o(t)\in\mathcal{C},\quad\text{as }t\downarrow 0.$ (5.6)
Equivalently, the distance between $x+tv$ and $\mathcal{C}$ is $o(t)$ as $t\downarrow 0$.
The set of vectors forward-tangent to $\mathscr{S}$ at $x$ in this sense is a closed cone, called the tangent cone to $\mathscr{S}$ at $x$. The set of vectors $v$ such that both $v$ and v-v are forward-tangent to $\mathscr{S}$ at $x$ is the tangent space at $x$. $\dashv$
Remark 5.1.
Fancier ideas of tangent cone are available[16, 17], but this simple one suffices for our needs.
For a smooth manifold, tangent spaces are the same as the tangent cones, so we do not need the latter concept. Consider, however, a pointed cone $\mathcal{C}$. If $v\in\mathcal{C}$, then $v$ is forward-tangent to $\mathcal{C}$ at 0, but v-v is not. If $\mathcal{C}$ is convex, then its tangent cones are also convex, and if $\mathcal{C}$ is a convex cone, the all its tangent cones are closed convex cones containing $\mathcal{C}$ itself. [Proof: for $x,v\colon\mathcal{C}$ and $0\geq t$, $x+tv\in\mathcal{C}$.]
Only vectors tangent to $\mathsf{CP}(\mathcal{H})$ at the identity are of much interest to us, so we introduce some appropriate notation for that case, deliberately modeled after a notational convention for Lie groups.
Definition 5.2 ($\mathsf{cp}^{+}(\mathcal{H})$$\mathsf{cp}(\mathcal{H})$).
The tangent cone to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ is denoted by $\mathsf{cp}^{+}(\mathcal{H})$, and the tangent space by $\mathsf{cp}(\mathcal{H})$. We may use $\mathsf{cp}^{+}$’ or $\mathsf{cp}$’ as an adjective, just like $\mathsf{CP}$’. $\dashv$
A CP curve $\sigma\colon[0,\epsilon[\rightarrow\mathsf{CP}(\mathcal{H})$, differentiable for $t>0$ and right differentiable at $t=0$, will be called smooth. If such a curve has $\sigma(0)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$, then the (right-) derivative at zero is in $\mathsf{cp}^{+}(\mathcal{H})$. It is not obvious that every operator in $\mathsf{cp}^{+}(\mathcal{H})$ can be “drawn out” into a smooth curve; eventually we will prove that.
Summing up the preceding discussion with the new notation, a continuously differentiable Markovian evolution family determines a continuous curve $\Lambda(t)$ in $\mathsf{cp}^{+}(\mathcal{H})$ according to (5.4). Suppose, conversely, that for any continuous curve $\Lambda(t)$ in $\mathsf{cp}^{+}(\mathcal{H})$, (5.4) with initial condition $\mathcal{E}(0,0)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ integrates to a curve in $\mathsf{CP}(\mathcal{H})$. (We will show this.) Then, for every such curve $\Lambda(t)$ we obtain a markovian evolution family because (5.5) also integrates to a curve in $\mathsf{CP}(\mathcal{H})$.
5.3 $L$-parametrization and the $L$-cone
Consider the map
$\begin{array}[]{ccccccl}\mathsf{HP}({\mathcal{H}})&\times&\mathrm{SA}(\mathcal{H})&\times&\mathrm{SA}(\mathcal{H})&\xrightarrow{\gamma}&\mathsf{HP}({\mathcal{H}})\\ (\Psi&,&G&,&H)&\mapsto&\Psi-{\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}}+G+iH).\end{array}$ (5.7)
Since the $\mathbb{R}$}-linear differential of $\bm{\theta}$} at $\operatorname{\mathbbm{1}}_{\mathcal{H}}$ is
$d{\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})\cdot(G+iH)=(G+iH)\,\square\,\operatorname{\mathbbm{1}}_{\mathcal{H}}+\operatorname{\mathbbm{1}}_{\mathcal{H}}\,\square\,(G-iH)=[G,\,\cdot\;]_{+}+[iH,\,\cdot\;],$ (5.8)
The differential $L{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}d\gamma(0)$ of $\gamma$ is
$\begin{array}[]{ccccccl}\mathsf{HP}({\mathcal{H}})&\times&\mathrm{SA}(\mathcal{H})&\times&\mathrm{SA}(\mathcal{H})&\xrightarrow{L}&\mathsf{HP}({\mathcal{H}})\\ (\Psi&,&G&,&H)&\mapsto&\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;].\end{array}$ (5.9)
Supposing $\Psi\in\mathsf{CP}(\mathcal{H})$, evidently $\gamma(\Psi,G,H)\in\mathsf{CP}(\mathcal{H})$. Therefore the curve $t\mapsto\gamma(t\Psi,tG,tH)$ is in $\mathsf{CP}(\mathcal{H})$ for $0\leq t$, so that $L(\Psi,G,H)\in\mathsf{cp}^{+}(\mathcal{H})$. Moreover, if $\Psi=0$, then the entire curve in $\mathsf{CP}(\mathcal{H})$, so $L(0,G,H)\in\mathsf{cp}(\mathcal{H})$.
Via $L$, a not-yet-determined subset of $\mathsf{cp}^{+}(\mathcal{H})$ is thus parametrized — nonuniquely — by the closed convex cone $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$. It, too, is a closed convex cone, which we will call the $L$-cone. Soon, we will see that $\mathsf{cp}^{+}(\mathcal{H})$ is the $L$-cone, so the terminology is redundant, but will be convenient in the interim.
Note that it is not true that only $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ is mapped to $\mathsf{cp}^{+}(\mathcal{H})$ by $L$. In fact, we can always subtract a large positive multiple of $\operatorname{\mathbbm{1}}_{\mathcal{H}}$ from $\Psi$ and compensate by shifting $G$ by half as much without changing the image under $L$.
Any triple $(\Psi,G,H)$ in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ such that $L(\Psi,G,H)=\Lambda$ is called an $L$-parametrization of $\Lambda$, and a restricted $L$-parametrization if, in case $\Lambda$ is $cp^{+}$, then $\Psi$ is CP. Significance of this restriction, and others, is considered in §5.8.
5.4 Integration for generators in the $L$-cone
Next, we show that a continuously-varying generator in the $L$-cone integrates to a Markovian CP evolution as in (5.5).
As discussed around (5.4), it is enough to show that the solution $\mathcal{E}$ of the initial value problem
$\dot{\mathcal{E}}(t)=\Lambda(t)\mathcal{E}(t),\quad\mathcal{E}(0)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ (5.10)
is CP. First, consider the special case of constant $\Lambda(t)\equiv\Lambda$, with solution $\exp t\Lambda$. It is CP by use of an elementary identity, as follows.
Lemma 5.1.
$\exp L(Z)$ is CP.
Proof.
since $\mathsf{CP}(\mathcal{H})$ is a monoid, for any natural number $n$,
$\gamma\left(\tfrac{Z}{n}\right)^{n}=\left(\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}+\tfrac{L(Z)+o(1)}{n}\right)^{n}$ (5.11)
is CP. And, since $\mathsf{CP}(\mathcal{H})$ is closed, the $n\to\infty$ limit of the display is also CP. That limit is $\exp L(Z)$. ∎
Next, move to a family of piecewise-constant approximations to (5.10), given, for $\epsilon>0$, by
$\Lambda_{\epsilon}(t)=\Lambda(n\epsilon)\text{ for }n\epsilon\leq t<(n+1)\epsilon.$ (5.12)
The solution, $\mathcal{E}_{\epsilon}(t)$ is $\exp t\Lambda(0)$ for $0\leq t\leq\epsilon$; $\exp[(t-\epsilon)\Lambda(\epsilon)]\cdot\exp\epsilon\Lambda(0)$ for $\epsilon\leq t\leq 2\epsilon$, etc. For any $t$, $\mathcal{E}_{\epsilon}(t)$ is simply a composite of exponentials, hence CP, since $\mathsf{CP}(\mathcal{H})$ is a monoid. Finally, since $\mathsf{CP}(\mathcal{H})$ is topologically closed, $\mathcal{E}(t)$ is CP if $\lim_{t\downarrow 0}\mathcal{E}_{\epsilon}(t)=\mathcal{E}(t)$. For this, it is important that $\Lambda(t)$ is continuous, so that $\Lambda_{\epsilon}(t)$ is uniformly close to $\Lambda(t)$ over bounded intervals. An estimate of the convergence is secured by the following general result.
Lemma 5.2.
Let
$\dot{x}=T(t)x,\quad\dot{x}^{\prime}=T^{\prime}(t)x^{\prime}$ (5.13)
be two linear differential equations in a normed vector space with the same initial condition $x(0)=x^{\prime}(0)=x_{0}$, and assume bounds
$0\leq t\leq T\Rightarrow\quad\|T(t)\|,\|T^{\prime}(t)\|\leq M,\quad\|T(t)-T^{\prime}(t)\|\leq\epsilon$ (5.14)
Then,
$0\leq t\leq T\Rightarrow\quad\|x(t)-x^{\prime}(t)\|\leq\epsilon e^{Mt}\|x_{0}\|.$ (5.15)
Proof.
The difference $\Delta(t)=x(t)-x^{\prime}(t)$ obeys the equation $\dot{\Delta}(t)=T(t)\Delta(t)+f(t)$ with initial condition $\Delta(0)=0$ and forcing term $f(t)=(T(t)-T^{\prime}(t))x^{\prime}(t)$. Now, combine the two bounds
$\|\Delta(t)\|\leq\int_{0}^{t}e^{M(t-s)}\|f(s)\|ds,\quad\|f(s)\|\leq\epsilon e^{Ms}\|x_{0}\|.$ (5.16)
5.5 Decomposition of $\mathrm{SA}(\mathfrak{B}(\mathcal{H}))$
$\mathfrak{J}$ effects an isometry of $\mathsf{CP}(\mathcal{H})$ onto $\mathrm{Pos}(\mathfrak{B}(\mathcal{H}))$ with $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}={\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})\;\xleftrightarrow{\;\mathfrak{J}\;}\;{\bm{\pi}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})$. Thus, we can learn about $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ by studying the tangent cone and tangent space to $\mathrm{Pos}(\mathfrak{B}(\mathcal{H}))$ at ${\bm{\pi}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})$, which we denote simply $C^{+}$ and $C$, respectively. Since vectors in these classes are in the ambient space $\mathrm{SA}(\mathfrak{B}(\mathcal{H}))$, we begin with considerations applying generally to the latter.
With $\mathfrak{B}(\mathcal{H})^{0}$ denoting the subspace of trace-zero operators in $\mathfrak{B}(\mathcal{H})$, we have an orthogonal decomposition $\mathfrak{B}(\mathcal{H})={\mathbb{C}}\operatorname{\mathbbm{1}}_{\mathcal{H}}\oplus\mathfrak{B}(\mathcal{H})^{0}$. With respect to this, any operator in $\mathrm{SA}(\mathfrak{B}(\mathcal{H}))$ has a “block-matrix” representation
$\left(\begin{array}[]{c|c}c&\overline{B}\\ \hline\cr B&\Phi\end{array}\right),\quad\text{with }c\in{\mathbb{R}},\;B\in\mathfrak{B}(\mathcal{H})^{0},\;\Phi\in\mathrm{SA}(\mathfrak{B}(\mathcal{H})^{0}).$ (5.17)
Flattened into an operator form, this is $\Phi+c{\bm{\pi}}[\operatorname{\mathbbm{1}}_{\mathcal{H}}]+B\overline{\operatorname{\mathbbm{1}}_{\mathcal{H}}}+\operatorname{\mathbbm{1}}_{\mathcal{H}}\overline{B}$, or, with the definition $K{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}B+\tfrac{c}{2}\operatorname{\mathbbm{1}}_{\mathcal{H}}$, $\Phi-K\overline{\operatorname{\mathbbm{1}}_{\mathcal{H}}}-\operatorname{\mathbbm{1}}_{\mathcal{H}}\overline{K}$. Here, in being implicitly promoted to an operator on $\mathfrak{B}(\mathcal{H})$, $\Phi$ is understood to annihilate the orthogonal complement (namely ${\mathbb{C}}\operatorname{\mathbbm{1}}_{\mathcal{H}}$) of its original domain $\mathfrak{B}(\mathcal{H})^{0}$. Splitting $K$ into hermitian and antihermitian parts as
$K=G+iH,$ (5.18)
$H$ must be trace-zero. Thus, we arrive at a linear isomorphism
$\begin{array}[]{ccccccl}\mathrm{SA}(\mathfrak{B}(\mathcal{H})^{0})&\times&\mathrm{SA}(\mathcal{H})&\times&\mathrm{SA}(\mathcal{H})^{0}&\xrightarrow[\cong]{\ell}&\mathrm{SA}(\mathfrak{B}(\mathcal{H}))\\ (\Phi&,&G&,&H)&\mapsto&\Phi-K\overline{\operatorname{\mathbbm{1}}_{\mathcal{H}}}-\operatorname{\mathbbm{1}}_{\mathcal{H}}\overline{K}.\end{array}$ (5.19)
The resemblance of this to $L$ prompts a definition analogous to $\gamma$ in (5.7):
$\begin{array}[]{ccccccl}\mathrm{SA}(\mathfrak{B}(\mathcal{H})^{0})&\times&\mathrm{SA}(\mathcal{H})&\times&\mathrm{SA}(\mathcal{H})^{0}&\xrightarrow{\sigma}&\mathrm{SA}(\mathfrak{B}(\mathcal{H}))\\ (\Phi&,&G&,&H)&\mapsto&\Phi-{\bm{\pi}}(\operatorname{\mathbbm{1}}_{\mathcal{H}}+G+iH).\end{array}$ (5.20)
Because $d{\bm{\pi}}(\operatorname{\mathbbm{1}}_{\mathcal{H}})\cdot K=K\overline{\operatorname{\mathbbm{1}}_{\mathcal{H}}}+\operatorname{\mathbbm{1}}_{\mathcal{H}}\overline{K}$, $\ell=d\sigma(0)$. And, since $\ell$ is an isomorphism, this says that for any any $\Lambda\in\mathrm{SA}(\mathfrak{B}(\mathcal{H}))$, there is a unique $(\Phi,G,H)\in\mathrm{SA}(\mathfrak{B}(\mathcal{H})^{0})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})^{0}$ such that $\Lambda$ is the derivative at zero of the curve $t\mapsto\sigma(t\Phi,tG,tH)$. Moreover, the half-curve $\{\sigma(t\Phi,tG,tH)\,:\,t>0\}$ is in $\mathrm{Pos}(\mathfrak{B}(\mathcal{H}))$ precisely when $\Phi$ is so, and the full curve, precisely when $\Phi=0$.
5.6 The $L$-cone is $\mathsf{cp}^{+}(\mathcal{H})$
The main conclusion of the preceding subsection is translated to $\mathsf{HP}(\mathcal{H})$ by means of
$\gamma(\mathfrak{J}\Phi,G,H)=\mathfrak{J}(\sigma(\Phi,G,H)).$ (5.21)
The statement is: $L$ effects an isomorphism $\mathsf{CP}^{0}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})^{0}\xrightarrow{\cong}\mathsf{cp}^{+}(\mathcal{H})$, where $\mathsf{CP}^{0}(\mathcal{H}){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\mathfrak{J}[\mathrm{Pos}(\mathfrak{B}(\mathcal{H})^{0})]$. In particular, $\mathsf{cp}^{+}(\mathcal{H})$ is contained in the $L$-cone, as was foreshadowed. With this, we have established that these classes of superoperators are the same:
  • generators of Markovian CP evolutions,
  • the $L$-cone,
  • $\mathsf{cp}^{+}(\mathcal{H})$.
Equivalence of the first two is a form of the GKSL theorem.
5.7 Lindblad parametrizers
5.7.1 Removing a redundancy
One part of the nonuniqueness of $L$-parametrization is trivial and should simply be removed. This is the insensitivity of $iH,\;\cdot\;]$ to shifting $H$ by a multiple of $\operatorname{\mathbbm{1}}_{\mathcal{H}}$. The way we deal with this is to replace the $\mathrm{SA}(\mathcal{H})$ summand corresponding to $H$ with the quotient space $\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. Alternatively, we could use $\mathrm{SA}(\mathcal{H})^{0}$ as was done in the previous subsection. Since $\mathrm{SA}(\mathcal{H})^{0}$ is a cross section of $\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$, the are essentially equivalent choices and formulas can be translated easily. However, the quotient space has the advantage of being applicable also to infinite-dimensional spaces.
5.7.2 The concept
The previous subsection showed that $\mathsf{cp}^{+}(\mathcal{H})$ is nonuniquely parametrized by $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$, via $L$. $L$ constructs a cp operator from a parametrization. How do we do the reverse — get a parametrization out?
Definition 5.3 (Lindblad parametrizer).
A Lindblad parametrizer (or just parametrizer) is a linear map $\Delta\colon\mathsf{HP}(\mathcal{H})\rightarrow$ $\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ which (i) is right-inverse to $L$, i.e., $L\circ\Delta$ is identity, and (ii) maps the closed convex cone $\mathsf{cp}^{+}(\mathcal{H})$ into the closed convex cone $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. $\dashv$
Remark 5.2.
Clause (ii) is the only reason this definition amounts to anything. For example, the trivial map $\Lambda\mapsto(\Lambda,0,0)$ is right-inverse to $L$. The point of this condition is that it requires $\Delta$ to pick out an $L$-parametrization of the kind referred to in the GKSL theorem when that is possible, namely one with CP $\Psi$.
We will see that the condition has surprising force, reducing the freedom to a choice of a subspace $\mathscr{M}$ complementary to $\mathsf{cp}(\mathcal{H})$ ($\mathsf{HP}(\mathcal{H})=\mathsf{cp}(\mathcal{H})\oplus\mathscr{M}$) satisfying $\mathscr{M}\cap\mathsf{cp}^{+}(\mathcal{H})\subseteq\mathsf{CP}(\mathcal{H})$.
For convenience, we introduce the abbreviation
$\mathsf{Lind\,{\mathcal{H}}}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}},$ (5.22)
which will be used from time to time.
Example 5.1 (Minimal Lindblad parametrizer, $\Delta^{\scriptstyle{min}}$).
Let us expose the minimal parametrizer we have already found in an explicit form, bypassing the Jamiołkowski transformation. Let $A_{0}=\tfrac{1}{\sqrt{\dim\mathcal{H}}}\operatorname{\mathbbm{1}}_{\mathcal{H}},A_{1},\ldots,A_{(\dim\mathcal{H})^{2}-1}$ be an orthogonal basis of $\mathfrak{B}(\mathcal{H})$. Then, $\Lambda\colon\mathsf{HP}(\mathcal{H})$ is decomposed on the basis $A_{i}\,\square\,A_{j}^{\dagger}$ (Example 2.1) as
$cA_{0}\,\square\,A_{0}+\sum_{i\geq 1}\left(d_{i}A_{i}\,\square\,A_{0}+\overline{d_{i}}A_{0}\,\square\,A_{i}^{\dagger}\right)+\sum_{i,j\geq 1}h_{ij}A_{i}\,\square\,A_{j}^{\dagger},$ (5.23)
where $h$ is a hermitian matrix. Reduce it to the form $h=\sum_{\alpha}\epsilon_{\alpha}b_{\alpha}b_{\alpha}^{\dagger}$ with the $b_{\alpha}$’s orthogonal column vectors and $\epsilon_{\alpha}=\pm 1$. Then, with the definitions $B_{\alpha}=\sum(b_{\alpha})_{i}A_{i}$ and $K=\tfrac{c}{2\dim\mathcal{H}}\operatorname{\mathbbm{1}}_{\mathcal{H}}+\tfrac{1}{\sqrt{\dim\mathcal{H}}}\sum d_{i}A_{i}$,
$\Lambda=\sum_{\alpha}\epsilon_{\alpha}{\bm{\theta}}(B_{\alpha})-K\,\square\,\operatorname{\mathbbm{1}}_{\mathcal{H}}-\operatorname{\mathbbm{1}}_{\mathcal{H}}\,\square\,K^{\dagger}.$ (5.24)
Now, with $\Psi=\sum\epsilon_{\alpha}{\bm{\theta}}(B_{\alpha})$ and K=GiHK=-G-iH, $\Lambda=L(\Psi,G,H)$. $\Lambda$ is in $\mathsf{cp}^{+}(\mathcal{H})$ iff $\Psi$ is CP, iff every $\epsilon_{\alpha}$ is +1+1. $\dashv$
5.7.3 Isomorphism between $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ and $\mathsf{cp}(\mathcal{H})$
[…]
$\mathsf{HP}(\mathcal{H})$$\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$$\mathsf{HP}(\mathcal{H})$$\mathsf{cp}^{+}(\mathcal{H})$$\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$$\mathsf{cp}^{+}(\mathcal{H})$$\mathsf{cp}(\mathcal{H})$$0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$$\mathsf{cp}(\mathcal{H})$$\scriptstyle{\Delta}$$\scriptstyle{L}$$\scriptstyle{\Delta}\scriptstyle{\subset}$$\scriptstyle{L}\scriptstyle{\subset}\scriptstyle{\subset}$$\scriptstyle{\Delta}\scriptstyle{\cong}\scriptstyle{\subset}\scriptstyle{\subset}$$\scriptstyle{L}\scriptstyle{\cong}\scriptstyle{\subset}$ (5.25)
We proceed to justify this.
Lemma 5.3.
The restriction of $L$ to $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ is an isomorphism onto $\mathsf{cp}(\mathcal{H})$. The restriction to $\mathsf{cp}(\mathcal{H})$ of any Lindblad parametrizer is thereby determined.
Proof.
$L$ is injective on $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$: Suppose $\Lambda{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}[G,\;\cdot\;]_{+}+[iH,\;\cdot\;]=0$. $2G=\Lambda\operatorname{\mathbbm{1}}_{\mathcal{H}}$, forcing $G=0$. And, the only operators commuting with all of $\mathfrak{B}(\mathcal{H})$ are multiples of the identity, so $H\in{\mathbb{R}}\operatorname{\mathbbm{1}}_{\mathcal{H}}$.
Now, let $\Delta$ be a parametrizer ($\Delta^{\scriptstyle{min}}$ is one!). By definition, it maps $\mathsf{cp}(\mathcal{H})$ into the convex cone $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. But, $\mathsf{cp}(\mathcal{H})$ is a vector space, hence must be mapped into the unique maximal vector space in that cone, namely $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$.
So, $\Delta$ and $L$ are both injections between the spaces indicated in the bottom row of (5.25). Since $L\circ\Delta$ is identity, they are therefore isomorphisms by dimension counting. ∎
It is good to recall again at this point that, necessarily, if the first component of $\Delta\Lambda$ is CP, then $\Lambda$ is $\mathsf{cp}^{+}$.
5.7.4 Structure of $\mathsf{cp}^{+}(\mathcal{H})$
Since $\Delta$ resticts to a fixed isomorphism of $\mathsf{cp}(\mathcal{H})$ onto $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$, it is determined by its restriction to a complementary subspace $\mathscr{M}$ (i.e., $\mathsf{HP}(\mathcal{H})=\mathsf{cp}(\mathcal{H})\oplus\mathscr{M}$) which is mapped into $\mathsf{HP}(\mathcal{H})\times 0\times 0$. However, $L(\Psi,0,0)=\Psi$, so the restriction of $\Delta$ to $\mathscr{M}$ must be simply $\Psi\mapsto(\Psi,0,0)$. Therefore, if $\Lambda=\Gamma+L(0,G,H)$ is a decomposition according to $\mathscr{M}\oplus\mathsf{cp}(\mathcal{H})$, then $\Delta\Lambda=(\Gamma,G,H)$.
However, not just any complementary subspace $\mathscr{M}$ is acceptable. According to clause (ii) of Def. 5.3, it must be such that
$\mathsf{cp}^{+}(\mathcal{H})\cap\mathscr{M}\subseteq\mathsf{CP}(\mathcal{H}).$ (5.26)
The minimal parametrizer $\Delta^{\scriptstyle{min}}$ corresponds to the choice $\mathscr{M}=\mathsf{cp}(\mathcal{H})^{\perp}$.
According to (5.26), the existence of parametrizers shows that
${\mathsf{cp}^{+}(\mathcal{H})}={\mathsf{cp}(\mathcal{H})}\,+\,\mathsf{CP}(\mathcal{H}).$ (5.27)
The two sides here are equal as sets. It was clear all along that both $\mathsf{cp}(\mathcal{H})$ and $\mathsf{CP}(\mathcal{H})$ are trivially subsets of $\mathsf{cp}^{+}(\mathcal{H})$. Now we see that it is the smallest convex cone containing them.
5.8 Significance of parametrizers
In the can-be-written formulation of the GKSL theorem, existence if a restricted $L$-parametrization appears merely as a test for membership in $\mathsf{cp}^{+}$. In practice, however, they (or even Kraus-expanded forms, see §5.9.2) are the preferred form in which to work with such operators. It would be considered a mistake to work with an $L$-parametrization $(\Psi,G,H)$ of a $cp^{+}$ map, for which $\Psi$ was not CP. Furthermore, for a $\mathsf{cp}$ map, one would insist on a parametrization with $\Psi$ equal to zero, not just CP. This very sensible restriction is not even hinted at in the can-be-written form. However, it comes out of the notion of Lindblad parametrizer automatically. These observations point to one significance of Lindblad parametrizers; they are a systematic way of obtaining good $L$-parametrizations.
But what are the physical ideas here? It is quite common to bundle $G$ and $H$ together into an effective Hamiltonian, leaving $\Psi$ as the “jump part” of a generator. For example, this idea underlies the integrability proof in Ref. [5]. Generators in $\mathsf{cp}(\mathcal{H})$ generate reversible evolutions and have unique $(0,G,H)$ parametrizations (remember that $H$ is an equivalence class). Reversibility is an unambiguous notion. A splitting into a reversible $G,H$ part and an irreversible $\Psi$ part is not; each choice of a parametrizer represents a distinct way of making that split. However, although not absolute, this notion of “irreversible part” is useful if used with caution.
Finally, in cases where one deals with continuously-varying $\mathsf{cp}^{+}$ maps, such as with time-variation or perturbation problems, one undoubtedly wants to know that continuous restricted $L$-parametrizations exist and how to get them. Lindblad parametrizations are an excellent answer to this need.
5.9 Trace-related properties and recovery of a conventional form
5.9.1 trace-non-increasing and trace-preserving
Turning very briefly to trace-related properties of Markovian CP evolution, note that
$\tfrac{d}{dt}\operatorname{\mathrm{Tr}}(\mathcal{E}(t,s)\rho)\Big|_{t=s}=\tfrac{d}{dt}\left\langle{\operatorname{\mathbbm{1}}_{\mathcal{H}}}\middle|{\mathcal{E}(t,s)\rho}\right\rangle\Big|_{t=s}=\left\langle{\Lambda(t)^{\dagger}\operatorname{\mathbbm{1}}_{\mathcal{H}}}\middle|{\rho}\right\rangle.$ (5.28)
Since $H,\operatorname{\mathbbm{1}}_{\mathcal{H}}]=0$, this leads to the constraints
$\begin{array}[]{llll}&\text{trace non-increasing:}&\Psi^{\dagger}\operatorname{\mathbbm{1}}_{\mathcal{H}}&\leq 2G\\ &\text{trace-preserving:}&\Psi^{\dagger}\operatorname{\mathbbm{1}}_{\mathcal{H}}&=2G.\end{array}$ (5.29)
This shows that any $\Lambda(t)$ can be modified trivially to yield a trace-non-increasing evolution by adding a multiple of the identity to $G$. Alternatively, and a little more sharply, since $\|e^{t\Lambda}\|\leq e^{t\|\Lambda\|}$, the modification $\Lambda\mapsto\Lambda-\|\Lambda\|\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$ serves.
5.9.2 Diagonal form
We recast a trace-preserving $L$-cone operator $\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$ in the conventional diagonal form (1.3). First, let $\Psi=\sum{\bm{\theta}}(A_{i})$ be a Kraus decomposition of $\Psi$. As pointed out in §4.1, these may be chosen mutually orthogonal. Then, applying the trace-preservation condition (5.29), $G=\tfrac{1}{2}\sum{\bm{\theta}}(A_{i})^{\dagger}\operatorname{\mathbbm{1}}_{\mathcal{H}}=\tfrac{1}{2}\sum A_{i}^{\dagger}A_{i}$. Combining corresponding terms, our operator is rewritten as
$\sum_{i}\left({\bm{\theta}}(A_{i})-\tfrac{1}{2}[A_{i}^{\dagger}A_{i},\;\cdot\;]_{+}\right)-[iH\;\cdot\;],$ (5.30)
which is (1.3) up to a small change of notation.
6 Recapitulation
  • $\mathsf{CP}$ is the largest subcategory of $\mathsf{Mon}$ compatible with tensor product and $\bm{\theta}$} is a monoidal functor.
  • Together with its distributivity over $\otimes$, almost everything we need to know about the Jamiołkowski transform $\mathfrak{J}$ is contained in the central diagram (3.12. It shows us that Kraus decomposition is extremal decomposition for the closed convex cone $\mathsf{CP}(\mathcal{H},\mathcal{K})$ and that $\bm{\theta}$} maps into the extreme CP maps.
  • Differentiable Markovian CP evolutions must satisfy the equation $\tfrac{\partial}{\partial t}\mathcal{E}(t,s)=\Lambda(t)\mathcal{E}(t,s)$ for $t\geq s$, and $\mathcal{E}(s,s)=\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$, with a generator $\Lambda(t)$ in $\mathsf{cp}^{+}(\mathcal{H})$, the tangent cone to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$. In fact, it is sufficient that $\Lambda(t)$ vary continuously in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$.
  • The tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ is parametrised nonuniquely via the linear $L$ map by $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. $L$ additionally maps the maximal subspace $\{0\}\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ isomorphically onto the maximal subspace in $\mathsf{cp}^{+}(\mathcal{H})$, namely the tangent space $\mathsf{cp}(\mathcal{H})$.
  • Lindblad parametrizers, right-inverses to $L$ that respect the cone structure by mapping $\mathsf{cp}^{+}(\mathcal{H})$ into $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$, and therefore $\mathsf{cp}(\mathcal{H})$ onto $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$, exist. $\Delta^{\scriptstyle{min}}$ is one.
7 Preparing for the separable setting
Starting in this section, we pivot toward the infinite-dimensional setting, more accurately, the separable setting. A separable Hilbert space is one which has a countable complete orthonormal system. More generally (and this applies equally to spaces like $\mathfrak{B}(\mathcal{H})$), a separable space is one which contains a countable dense subset. Non-separable spaces are not needed for physics.
7.1 Trace-class/bounded duality
This subsection does not treat its subject in a strictly logical order. For more systematic developments, see textbook presentations in Chapter 3 of Ref. [18], §3.4 of Ref. [19], or Chapter 7 exercises of Ref. [20].
Since Hilbert space structures on all our spaces played such a central rôle in the finite-dimensional theory, e.g., the Jamiołkowski transform was a unitary map, we will approach the separable setting by seeing how that falls apart. There is a space of linear operators on $\mathcal{H}$ — the Hilbert-Schmidt operators — for which a Hilbert space structure can be defined just as before. But the associated topology is at the same time too strong and too weak. We wind up with a rigged Hilbert space
$\mathcal{B}^{1}(\mathcal{H})\subset\mathcal{B}^{2}(\mathcal{H})\subset\mathfrak{B}(\mathcal{H}).$ (7.1)
$\mathfrak{B}(\mathcal{H})$ is the space of bounded operators on $\mathcal{H}$. We have changed notation from $\mathfrak{B}(\mathcal{H})$ to emphasize the topologies on these spaces. A positive operator $A$ has a trace in $0,\infty]$ unambiguously defined by
$\operatorname{\mathrm{Tr}}A=\sum_{i}\left\langle{h_{i}}\middle|{Ah_{i}}\right\rangle,$ (7.2)
where $\{h_{i}\}$ is any complete orthonormal system in $\mathcal{H}$. For such an operator, the trace “norm” $\|A\|_{1}$ is equal to its trace, This definition extends to a general bounded operator $B$ by use of polar decomposition: if $B=MU$ with $M\geq 0$ and $U$ unitary, then $\|B\|_{1}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\|M\|_{1}$. If $\|A\|_{1}<\infty$, then $A$ is said to be trace-class. In that case, its trace is unambiguously given by (7.2). $\mathcal{B}^{1}(\mathcal{H})$, the space of all trace-class operators, is a Banach space under this norm. $A$ is Hilbert-Schmidt if $A^{\dagger}A$ is trace class. On the Hilbert-Schmidt operators we have the by-now familiar inner product $\left\langle{A}\middle|{B}\right\rangle=\operatorname{\mathrm{Tr}}A^{\dagger}B$, and norm $\|A\|_{2}=\sqrt{\left\langle{A}\middle|{A}\right\rangle}$. $\mathfrak{B}(\mathcal{H})$ is the space of all bounded linear operators, with operator norm $\|A\|=\sup\{\|Ah\|\,:\,\|h\|=1\}$. It, too, is a Banach space. The Hilbert-Schmidt inner product $\left\langle{A}\middle|{B}\right\rangle$ does not extend to $\mathfrak{B}(\mathcal{H})$, but we can allow $A$ to be any bounded operator if we simultaneously restrict $B$ to be trace-class. This pairing identifies $\mathfrak{B}(\mathcal{H})$ with the Banach space dual of $\mathcal{B}^{1}(\mathcal{H})$ as $B\mapsto\left\langle{B}\middle|{\;\cdot\;}\right\rangle$. However, $\mathcal{B}^{1}(\mathcal{H})$ is not the dual space of $\mathfrak{B}(\mathcal{H})$, i.e., $\mathcal{B}^{1}(\mathcal{H})$ is not reflexive. That $\mathfrak{B}(\mathcal{H})$ is the dual space of $\mathcal{B}^{1}(\mathcal{H})$ means that the operator norm satisfies
$\|B\|=\sup\left\{|\left\langle{B}\middle|{\rho}\right\rangle|\,\colon\,\|\rho\|_{1}\leq 1\right\}.$ (7.3)
Conversely,
$\|\rho\|_{1}=\sup\left\{|\left\langle{B}\middle|{\rho}\right\rangle|\,\colon\,\|B\|\leq 1\right\}.$ (7.4)
Actually, in this last formula, it is enough to consider finite-rank operators $B$. Here is another fact about the interplay of these spaces which will be useful in the following. Since $BA$ runs over some subset of the ball of radius $\|A\|$ as $B$ runs over the unit ball, $\langle B,A\rho\rangle=\langle BA,\rho\rangle$ implies that $\|A\rho\|_{1}\leq\|A\|\|\rho\|_{1}$.
The superoperators which interest us are natively in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. Such a $\Lambda\in\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$ has adjoint $\Lambda^{\dagger}\in\mathfrak{B}(\mathfrak{B}(\mathcal{H}))$ defined by
$\left\langle{\Lambda^{\dagger}B}\middle|{\rho}\right\rangle=\left\langle{B}\middle|{\Lambda\rho}\right\rangle.$ (7.5)
The norm on $\Lambda$ is just the operator norm, as for an operator on any Banach space:
$\|\Lambda\|=\sup\left\{\|\Lambda\rho\|_{1}\,\colon\,\|\rho\|_{1}\leq 1\right\}.$ (7.6)
Remark 7.1.
Usually, one uses a pairing $\mathfrak{B}(\mathcal{H})\times\mathcal{B}^{1}(\mathcal{H})$ which is linear in both variables: $(B,\rho)\mapsto\langle B,\rho\rangle=\operatorname{\mathrm{Tr}}B\rho$. Then, one defines the dual of a superoperator as $\langle\Lambda^{*},\rho\rangle=\langle B,\Lambda\rho\rangle$. For a general Banach space in place of $\mathcal{B}^{1}(\mathcal{H})$, this would be the only choice. We are using instead a sesquilinear pairing in keeping with the rigged Hilbert space idea and for maximum continuity with the finite-dimensional theory.
7.2 Strong-operator topology
The norm of a normed (e.g., Banach) space $(\mathscr{X},\|\cdot\|)$ defines a metric on that space, which allows the discussion of convergence of sequences, neighborhoods, and open and closed sets. The distance between $x$ and $y$ is $\|x-y\|$, and a sequence $(x_{n})$ converges to $x$ if $\|x-x_{n}\|\to 0$. Another way to describe this is in terms of neighborhoods of $x$. We write $\mathrm{ball}(\mathscr{X})$ for the centered open ball $\left\{x\colon\mathscr{X}\,\colon\,\|x\|<1\right\}$ of radius $1$, and $r\cdot\mathrm{ball}(\mathscr{X})$ for its dilation by a factor $r$. Then $x+r\cdot\mathrm{ball}(\mathscr{X})$ is the set of points less than distance $r$ from $x$, and a sequence converges to $x$ if it eventually stays in each such neighborhood of $x$. A set $S$ is a neighborhood of $x$ if it contains such a ball around $x$, and is open if it is a neighborhood of each of its points. Complements of open sets are closed. A principal significance of $C$ being closed is that, if a sequence within $C$ has a limit, it is also in $C$.
Now, if $\mathcal{X}$ and $\mathcal{Y}$ are both normed spaces, then $\mathfrak{B}(\mathcal{X},\mathcal{Y})$ is a normed space with the operator norm. But there are other natural norm-like functions. Consider $\mathscr{A}\mapsto\|\mathscr{A}x\|$ for some $x\colon\mathcal{X}$, which has all the properties of a norm except that $\|\mathscr{A}x\|=0$ does not guarantee that $\mathscr{A}=0$. Such an object is called a seminorm. We can use these to define open “strong balls” $\mathrm{ball}_{S}(x)=\left\{\mathscr{A}\,\colon\,\|\mathscr{A}x\|<1\right\}$, $\mathrm{ball}_{S}(x,y)=\mathrm{ball}_{S}(x)\cap\mathrm{ball}_{S}(y)$, and so forth for any finite set $x,\ldots,y$. These balls can be used to define the other notions as we did with the norm topology (i.e., they are basic neighborhoods of zero). For instance, sequence $\mathscr{A}_{n}$ converges to $\mathscr{A}$ in the strong-operator topology (SOT) means that $\|\mathscr{A}x-\mathscr{A}_{n}x\|\to 0$ for every $x\colon\mathcal{X}$.
Lemma 7.1 below collects some more-or-less-standard properties of SOT which will be useful in the sequel. In the proof, appeal is made to the Banach-Steinhaus theorem, which is a form of the Uniform Boundedness Principle and says: suppose $\mathscr{A}_{n}\colon\mathcal{X}\rightarrow\mathcal{Y}$ is a sequence of bounded operators from Banach space $\mathcal{X}$ to Banach space $\mathcal{Y}$. If, for every $x\in\mathcal{X}$, the sequence $\mathscr{A}_{n}x$ is convergent in $\mathcal{Y}$, then, in fact, the map $x\mapsto\lim\mathscr{A}_{n}x$ defines a bounded operator. Discussion of the Uniform Boundedness Principle can be found in most functional analysis texts, for instance §9.3 of Ref. [20], §2.2 of Ref. [19], §III.5 of Ref. [21], §III.14 of Ref. [22]. We also use the following simple idea. If $(x_{n})$ is a sequence in a Banach space, and for any $\epsilon>0$, there is a Cauchy sequence $(x_{n}^{\epsilon})$ with $\|x_{n}-x_{n}^{\epsilon}\|<\epsilon$ for all $n$, then the original sequence $(x_{n})$ is Cauchy. This will be referred to as the “shadowing lemma”.
Lemma 7.1.
Strong-operator convergence has the following properties.
  1. (a)
    $\mathscr{A}_{n}\xrightarrow{\text{SOT}}\mathscr{A}$, $\mathscr{B}_{n}\xrightarrow{\text{SOT}}\mathscr{B}$ imply $\mathscr{A}_{n}\mathscr{B}_{n}\xrightarrow{\text{SOT}}\mathscr{A}\mathscr{B}$.
  2. (b)
    $\mathscr{A}_{n}\xrightarrow{\text{SOT}}\mathscr{A}$ implies $\|\mathscr{A}\|\leq\liminf\|\mathscr{A}_{n}\|$.
  3. (c)
    $\{\|\mathscr{A}_{n}\|\}$ bounded and $(\mathscr{A}_{n}x)$ converges for all $x$ in some dense set implies that $(\mathscr{A}_{n})$ converges SOT.
  4. (d)
    $A_{n}\stackrel{{\scriptstyle SOT}}{{\longrightarrow}}A$ in $\mathfrak{B}(\mathcal{H})$ implies $A_{n},\cdot\;]_{\pm}\stackrel{{\scriptstyle SOT}}{{\longrightarrow}}[A,\cdot\;]_{\pm}$ in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$.
Proof.
(a). The hypotheses immediately imply that for any $x$, $\mathscr{A}x-\mathscr{A}_{n}x\to 0$ and $(\mathscr{B}-\mathscr{B}_{n})\mathscr{A}x\to 0$. Moreover, by the Uniform Boundedness Principle, $\{\|\mathscr{B}_{n}\|\}$ is bounded, so that $\mathscr{B}_{n}(\mathscr{A}x-\mathscr{A}_{n}x)\to 0$, as well. Thus,
$\mathscr{B}\mathscr{A}x-\mathscr{B}_{n}\mathscr{A}_{n}x=(\mathscr{B}-\mathscr{B}_{n})\mathscr{A}x+\mathscr{B}_{n}(\mathscr{A}x-\mathscr{A}_{n}x)\to 0.$ (7.7)
(b). Let $\|x\|=1$ be such that $\|\mathscr{A}x\|>\|\mathscr{A}\|-\epsilon$. Then, $\|\mathscr{A}_{n}x\|>\|\mathscr{A}\|-\epsilon$ for large enough $n$, so $\limsup\|\mathscr{A}_{n}\|\geq\|\mathscr{A}\|-\epsilon$.
(c). By the norm bound, $\|\mathscr{A}_{n}y^{\epsilon}-\mathscr{A}_{n}x\|<\epsilon$ for $y^{\epsilon}$ close enough to $x$, and there is such $y^{\epsilon}$ in the given dense set. Now apply the shadowing lemma.
(d). For any $v\in\mathcal{H}$, $A_{n}v\to Av$ implies that $A_{n}{\bm{\pi}}[v]\to A{\bm{\pi}}[v]$ in $\mathcal{B}^{1}(\mathcal{H})$ ($A_{n}{\bm{\pi}}[v]$ is just operator multiplication), and therefore $A_{n}\rho\to A\rho$ if $\rho\in\mathcal{B}^{1}(\mathcal{H})$ has finite rank. Since $\mathcal{B}^{\scriptstyle{fin}}{\mathcal{H}}$ is dense in $\mathcal{B}^{1}(\mathcal{H})$, the shadowing lemma implies $A_{n}\rho\to A\rho$ for any $\rho$. A similar argument shows $\rho A_{n}\to\rho A$. Now, just add or subtract. ∎
7.3 Weak-operator topology
Convergence in the weak-operator topology (WOT) is an even more permissive mode of convergence. A sequence $A_{n}\in\mathfrak{B}(\mathcal{H})$ converges in WOT if $\left\langle{w}\middle|{A_{n}v}\right\rangle$ converges for every $w,v\in\mathcal{H}$. By a double application of the Banach-Steinhaus theorem, the limits define a bounded operator. In this Hilbert space case, it is sufficient to check just diagonal matrix elements $\left\langle{v}\middle|{A_{n}v}\right\rangle$, as the others follow by polarization. The concept of WOT convergence also applies to superoperators $(\Gamma_{n})$ in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$ thus: $\left\langle{B}\middle|{\Gamma_{n}\rho}\right\rangle$ converges for every $B\in\mathfrak{B}(\mathcal{H})$, $\rho\in\mathcal{B}^{1}(\mathcal{H})$. Alternatively, we may discuss WOT in terms of subbasic neighborhoods, just as for SOT. $A\mapsto|\left\langle{w}\middle|{Av}\right\rangle|$ is a seminorm, corresponding to a “weak ball” $\mathrm{ball}_{W}(w,v)=\left\{A\in\mathfrak{B}(\mathcal{H})\,\colon\,|\left\langle{w}\middle|{Av}\right\rangle|<1\right\}$.
Norm convergence implies SOT convergence, which in turn implies WOT convergence.
7.4 Meanings of old notations in the separable setting
Much of the notation used in the finite-dimensional setting transfers smoothly. The source of most of the clarifications needed arise from the fact that many of the occurences of $\mathfrak{B}(\mathcal{H})$” should be replaced by $\mathcal{B}^{1}(\mathcal{H})$”. In particular, the superoperators which interest us are natively in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. Hence we have
$\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})\subset\mathsf{HP}(\mathcal{H},\mathcal{K})\subset\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{K})).$ (7.8)
This is a matter not just of which linear maps are in those spaces, but what the default norm is. Then, we have
$\mathsf{Lind\,{\mathcal{H}}}=\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}},$ (7.9)
just as before. $\mathrm{Pos}(\mathcal{H})$ is now ambiguous. We will use an appended “++”, i.e., $\mathcal{B}^{1}(\mathcal{H})^{+}$ and $\mathfrak{B}(\mathcal{H})^{+}$ are the positive elements in the indicated spaces.
8 Filtrations: sequences of finite-dimensional approximations
In the preceding sections, we have built up a very satisfactory theory of $\mathsf{CP}$ maps and $\mathsf{cp}^{+}$ maps over finite dimensional Hilbert spaces. The basic idea of the present approach is to make use of that theory through finite-dimensional “approximations” of various things defined over a separable Hilbert space.
8.1 Motivational sketch
Let $\mathcal{H}_{\alpha}$ be the span of $\varphi_{1},\varphi_{2},\ldots,\varphi_{\alpha}$, where $\varphi_{1},\varphi_{2},\ldots$ is a complete orthonormal system in $\mathcal{H}$. Given $h$ in $\mathcal{H}$, we can project it orthogonally into the $\mathcal{H}_{\alpha}$, obtaining a sequence $P_{1}h,P_{2}h,\ldots$ of approximations.
Now, the $\mathcal{H}_{\alpha}$ are (finite-dimensional) Hilbert spaces in their own right, as well as subspaces of $\mathcal{H}$. We can take an operator $A$ on $\mathcal{H}_{\alpha}$ and regard it as a finite-rank operator on $\mathcal{H}$. $Ah$ is computed by applying $A$ as given to $P_{\alpha}h$ and then regarding the resulting vector as a member of $\mathcal{H}$. Conversely, given an operator $A$ on $\mathcal{H}$, we get one on $\mathcal{H}_{\alpha}$, as a projection $\hat{P}_{\alpha}A$, given as $(\hat{P}_{\alpha}A)h=P_{\alpha}(Ah)$. Thus, we get $\mathfrak{B}(\mathcal{H}_{1}),\mathfrak{B}(\mathcal{H}_{2}),\ldots$ as nested subspaces of $\mathfrak{B}(\mathcal{H})$, and $\mathcal{B}^{1}(\mathcal{H}_{1}),\mathcal{B}^{1}(\mathcal{H}_{2}),\ldots$ as nested subspaces of $\mathcal{B}^{1}(\mathcal{H})$. $\mathfrak{B}(\mathcal{H}_{\alpha})$ and $\mathcal{B}^{1}(\mathcal{H}_{\alpha})$ contain exactly the same linear operators; it is just their norms that are different.
We can copy the pattern at the next level, obtaining $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}_{\alpha}))$ as a subspace of $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. This is really the level at which we want to apply these ideas. For, given $\Lambda$ in $\mathsf{HP}(\mathcal{B}^{1}(\mathcal{H}))$, we can project it into $\mathsf{HP}(\mathcal{B}^{1}(\mathcal{H}))$, apply the minimal Lindblad parametrizer to obtain a representation $Z_{\alpha}$ in $\mathsf{Lind\,{\mathcal{H}_{\alpha}}}$, and regard $Z_{\alpha}$ as an element of $\mathsf{Lind\,{\mathcal{H}}}$. Is it the case, perhaps, that the $Z_{\alpha}$ converge to a Lindblad representation of $\Lambda$? Not quite. It turns out that some subsequence does converge to a Lindblad parametrization of $\Lambda$, in an appropriate metric.
The rest of this section, and the rest, are devoted to setting up the machinery needed to fulfill that promise.
8.2 Formal development
The examples are the important thing in this subsection. It is not a great exaggeration to say that the rest is just there to support them.
Definition 8.1 (Prefiltration).
Let $\mathscr{X}$ be a Banach space, and $\mathscr{P}$ a sequence $(\mathscr{P}_{\alpha})$ in $\mathfrak{B}(\mathscr{X}))$ with ranges $(\mathscr{X}_{\alpha})$. $(\mathscr{X},\mathscr{P})$ is a prefiltration if
  1. (F1)
    $\{\|\mathscr{P}_{\alpha}\|\}$ is bounded
  2. (F2)
    $\forall\alpha$, $\operatorname{ran}\mathscr{P}_{\alpha}$ is finite-dimensional
  3. (F3)
    $\forall\alpha,\beta$, $\mathscr{P}_{\alpha}\mathscr{P}_{\beta}=\mathscr{P}_{\min(\alpha,\beta)}$
  4. (F4)
    $(\forall\alpha,\,\mathscr{P}_{\alpha}x=0)$ implies $x=0$
Paraphrasing, the $\mathscr{P}_{\alpha}$ comprise a uniformly bounded sequence of finite-dimensional, nested, and collectively injective projections.
$\dashv$
Definition 8.2 (Adapted sequence, projective sequence).
A sequence $(x_{\alpha}\colon\mathscr{X})$ is adapted to the prefiltration $(\mathscr{X},\mathscr{P})$ if $x_{\alpha}\in\mathscr{X}_{\alpha}$ for all $\alpha$. It is projective if, furthermore,
$\alpha\leq\beta\;\Rightarrow\;x_{\alpha}=\mathscr{P}_{\alpha}x_{\beta},$ (8.1)
and eventually projective if (8.1) holds for $\alpha$ exceeding some threshold value.
$\dashv$
Given a prefiltration $(\mathscr{X},\mathscr{P})$ and $x\colon\mathscr{X}$, the projections $x_{\alpha}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\mathscr{P}_{\alpha}x$ comprise a bounded projective sequence. By (F4), the sequence $(x_{\alpha})$ of projections determines $x$. Each $x$ corresponds to a distinct bounded projective sequence. But, we need to work the other way round, determining $x$ from its projections. Again by (F4), any bounded projective sequence $(x_{\alpha})$ is the projections of a unique $x$ if there is any such $x$ at all. But there is no guarantee of that. If there is, then the situation is nice: bounded projective sequences are in one-to-one correspondence with elements of $\mathscr{X}$.
Definition 8.3.
A prefiltration $(\mathscr{X},\mathscr{P})$ is a filtration if
  1. 5.
    Whenever $(x_{\alpha}\colon\mathscr{X}_{\alpha})$ is projective and bounded, there exists $x:\mathscr{X}$ such that $x_{\alpha}=\mathscr{P}_{\alpha}x$.
If this situation guarantees $x_{\alpha}\to x$ in norm, then $(\mathscr{X},\mathscr{P})$ is a norm filtration. If the filtration is on a function space and the guaranteed convergence is in SOT sense, it is a SOT filtration. $\dashv$
Example 8.1 ($(\mathcal{H},P)$ is a norm filtration).
Consider a Hilbert space $\mathcal{H}$ with complete orthonormal system $(\varphi_{k})$ and $\mathscr{P}_{\alpha}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}P_{\alpha}$ is the orthoprojection onto the span $\mathcal{H}_{\alpha}$ of $\{\varphi_{1},\ldots,\varphi_{\alpha}\}$. This is our prototype of a prefiltration. If we think of the corresponding column vector representation, then $\mathscr{P}_{\alpha}$ is the operation of zeroing all except for the first $\alpha$ entries. More generally, we may take $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ to be an increasing sequence of finite-dimensional subspaces which are collectively dense in $\mathcal{H}$, i.e., $\overline{\cup_{\alpha\in{\mathbb{N}}}\mathcal{H}_{\alpha}}=\mathcal{H}$.
To see that this is actually a norm filtration, suppose $(v_{\alpha})$ is projective and bounded. Then $v_{\alpha}=P_{\alpha}v_{\alpha+1}$, so $\|v_{n+1}\|^{2}=\|v_{n}\|^{2}+\|v_{n+1}-v_{n}\|^{2}$, so $v_{\alpha}=\sum_{\beta=1}^{\alpha}(v_{\beta}-v_{\beta-1})$ is a sum of orthogonal increments whose squared norms are summable. Hence $v_{\alpha}$ converges, to say $v$, and $P_{\alpha}v=v_{\alpha}$. For our prototype, the sequence $(v_{\alpha})$ corresponds to starting with a zero vector and setting the components to $\left\langle{\varphi_{\alpha}}\middle|{v}\right\rangle$ one-by-one. $\dashv$
Example 8.2.
For filtrations $(\mathscr{X},\mathscr{P})$ and $(\mathscr{X}^{\prime},\mathscr{P}^{\prime})$, the direct sum $(\mathscr{X}\times\mathscr{X}^{\prime},\mathscr{P}_{n}\times\mathscr{P}^{\prime})$ is another. $\dashv$
A new tool is needed to deal with the spaces $\mathfrak{B}(\mathcal{H},\mathcal{K})$ and $\mathcal{B}^{1}(\mathcal{H})$.
Lemma 8.1.
If $(\mathscr{X},\mathscr{P})$ and $(\mathscr{X}^{\prime},\mathscr{P}^{\prime})$ are norm filtrations, then $(\mathscr{P}^{\prime}_{\alpha}\,\square\,\mathscr{P}_{\alpha})$ is a SOT filtration on $\mathfrak{B}(\mathscr{X},\mathscr{X}^{\prime})$
Proof.
F1 – F3: Straightforward.
F4: Assume $\mathscr{P}^{\prime}_{\alpha}T\mathscr{P}_{\alpha}=0$ for all $\alpha$. Since $(\mathscr{X},\mathscr{P})$ is a norm filtration, for every $x$, $\mathscr{P}_{\alpha}x\to x$ and $T\mathscr{P}_{\alpha}x\to Tx$, and finally $\mathscr{P}^{\prime}_{\alpha}T\mathscr{P}_{\alpha}x\to\mathscr{P}^{\prime}_{\alpha}Tx$. So, $\mathscr{P}^{\prime}_{\alpha}(Tx)=0$ and by (F4) for $(\mathscr{X}^{\prime},\mathscr{P}^{\prime})$, $Tx=0$, i.e., $T=0$.
F5: Now, suppose $(T_{\alpha})$ is a bounded projective sequence, and take $x\in\mathscr{X}_{\alpha}$. Then, for $\alpha\leq\beta\leq\gamma$, $T_{\beta}x=\mathscr{P}_{\beta}T_{\gamma}\mathscr{P}_{\beta}x=\mathscr{P}_{m}(T_{\gamma}x)$. Thus, $(T_{\alpha}x\in\mathscr{X}^{\prime}_{\alpha})$ is bounded and eventually projective, and therefore converges in norm (by hypothesis). That is, $(T_{\alpha})$ is a bounded sequence converging on dense subset $\cup_{\alpha}\mathscr{X}_{\alpha}$ of $\mathscr{X}$ and therefore converging in SOT, so say $T$. Then, $\mathscr{P}^{\prime}_{\beta}T\mathscr{P}_{\beta}x=\mathscr{P}^{\prime}_{\beta}\lim_{\alpha\to\infty}T_{\alpha}\mathscr{P}_{\beta}x=\lim_{\alpha\to\infty}\mathscr{P}^{\prime}_{\beta}T_{\alpha}\mathscr{P}_{\beta}x=$ (projectivity) $\lim_{\alpha\to\infty}T_{\beta}x=T_{\beta}x$
Example 8.3 ($(\mathfrak{B}(\mathcal{H},\mathcal{K}),P^{\prime}\,\square\,P)$ is a SOT filtration).
With $(\mathcal{H},P)$ and $(\mathcal{K},P^{\prime})$ as in Example 8.1, this is just an application of the preceding Lemma.
For the matrix representation introduced in Example 8.1, $\mathscr{X}_{\alpha}$ is the operators which are nonzero except for the upper-left $\alpha\times\alpha$ block, and $P^{\prime}_{\alpha}\,\square\,P_{\alpha}$ amounts to zeroing all entries outside that block. $P^{\prime}_{\alpha}\mathfrak{B}(H)P_{\alpha}$ can be identified with $\mathfrak{B}(\mathcal{H}_{\alpha},\mathcal{K}_{\alpha})$ by throwing away everything except the $\alpha\times\alpha$ block. $\dashv$
Example 8.4 ($(\mathcal{B}^{1}(\mathcal{H}),{\bm{\theta}}(P))$ is a norm filtration).
That this is a prefiltration follows almost automatically from Example 8.3. We just need to check boundedness of the projections because the norm is different: $\|PAP\|_{1}\leq\|P\|\|AP\|_{1}\leq\|P\|\|A\|_{1}\|P\|$.
If $(A_{\alpha})$ is projective and bounded in trace-norm, then it is operator-norm bounded, so by the previous example $A_{\alpha}\xrightarrow{SOT}A$ and $A_{\alpha}=P_{\alpha}AP_{\alpha}$ for some $\in\mathfrak{B}(\mathcal{H}). AA$ is actually trace-class: Let $M$ be a bound for the trace norms of the $A_{\alpha}$. If $A$ is not trace class, there is $B\in\mathfrak{B}(\mathcal{H})$ with $\|B\|\leq 1$ such that $\left\langle{B}\middle|{A}\right\rangle>M$. But WOT convergence is weaker than SOT, so $A_{\alpha}\xrightarrow{WOT}A$, which implies that $\left\langle{B}\middle|{A_{\alpha}}\right\rangle>M$ for large enough $\alpha$, a contradiction. We conclude that $(\mathcal{B}^{1}(\mathcal{H}),{\bm{\theta}}(P))$ is a SOT filtration.
But we made the stronger claim that it is a norm filtration, and to justify that we need to show
(*) for $A\colon\mathcal{B}^{1}(\mathcal{H})$, $P_{\alpha}AP_{\alpha}\to A$ in trace norm.
Observe that $A\in\mathcal{B}^{1}(\mathcal{H})$ is approximated by a finite linear combination of rank-1 projectors ${\bm{\pi}}[v]$, and any such projector is approximated by some ${\bm{\pi}}[w]$ with $w\in\cup\mathcal{H}_{\alpha}$. [find $w\in\cup\mathcal{H}_{\alpha}$ an approximation to $v$, say $w=v-\delta$. Then, ${\bm{\pi}}[v]-{\bm{\pi}}[w]=\delta\overline{v}+v\overline{\delta}-{\bm{\pi}}[\delta]$ has trace norm less than $2\|v\|\|\delta\|+\|\delta\|^{2}$.] So, $\cup_{\alpha}\mathcal{B}^{1}(\mathcal{H}_{\alpha})$ is dense in $\mathcal{B}^{1}(\mathcal{H})$ in trace norm.
Now, returning to (*), note that $\|P_{\alpha}AP_{\alpha}\|_{1}\leq\|P_{\alpha}\|^{2}\|A\|_{1}$, so that by boundedness of $\{\|P_{\alpha}\|\}$, $\{{\bm{\theta}}(P_{\alpha})\}$ is uniformly bounded in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. Therefore, we only need to show (*) for $A$ in the dense set $\cup_{\alpha}\mathcal{B}^{1}(\mathcal{H}_{\alpha})$. But this is trivial: for $A\in\mathcal{B}^{1}(\mathcal{H}_{\alpha})$, $P_{\beta}AP_{\beta}=A$ as soon as $\alpha\leq\beta$. $\dashv$
Remark 8.1.
Example 8.3 involved two possibly distinct Hilbert spaces. Had there been only one, would it have been okay to write the projection as ${\bm{\theta}}(P_{\alpha})$, as in Example 8.4? We adopt the following point of view: $\square\,$ is just a syntactic device. It merely shows where the argument goes. But, ${\bm{\theta}}(A)$ specifically means an operator in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H},\mathcal{K}))$, not an operator in $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$. The latter, if it were needed, would be denoted ${\bm{\theta}}(A)^{\dagger}$”.
Example 8.5 ($(\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H})),{\bm{\theta}}(P)\,\square\,{\bm{\theta}}(P))$ is a SOT filtration).
This is a straightforward application of Lemma 8.1 to example 8.4. $\dashv$
Example 8.6 ($\mathsf{Lind\,{\mathcal{H}}}$ is a SOT filtration).
($\mathsf{HP}(\mathcal{H})$ is a real Banach subspace of $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$, and $\mathrm{SA}(\mathcal{H})$ of $\mathfrak{B}(\mathcal{H})$. So $\mathsf{Lind\,{\mathcal{H}}}=\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ is a SOT filtration under the product of the projections on its factors, because each of those is a SOT filtration according to preceding examples. $\dashv$
8.3 Generic notation
Unpacked (partially), the projections on $\mathsf{Lind\,{\mathcal{H}}}$ in Example 8.6 are
$({\bm{\theta}}(P_{\alpha})\,\square\,{\bm{\theta}}(P_{\alpha}))\times(P_{\alpha}\,\square\,P_{\alpha})\times(P_{\alpha}\,\square\,P_{\alpha}).$ (8.2)
It is completely impracticable to be using that expression. However, it is in some sense not really that complicated, as it arises by fairly straightforward constructions starting from $P_{\alpha}$. Since all of the projections we will be using in §§10 and 11 arise from simple projections $P_{\alpha}$ on one (maybe two) Hilbert space in just the way laid out in the Examples, we will mostly use the unified abstract notation $\mathscr{P}_{\alpha}$ for all of them. Thus, context matters for proper parsing and ambiguity is possible. If using this notation, we cannot write simply $\mathscr{P}_{\alpha}\Lambda\rho$”. We must write $(\mathscr{P}_{\alpha}\Lambda)\rho$” or $\mathscr{P}_{\alpha}(\Lambda)\rho$”, where the projection acts on $\Lambda$; or $\mathscr{P}_{\alpha}(\Lambda\rho)$”, where it acts on $\Lambda\rho$.
9 A filtration induces a complete, boundedly-precompact metric on its base space
Recall that the fundamental characteristic of a filtration $(\mathscr{X},\mathscr{P})$ is that $\mathscr{X}$ is in one-to-one correspondence with the bounded projective sequences. We can use this to transfer structure, such as a metric, on the latter to the former. The metric, $d$, we define on $\mathscr{X}$ this way has the very useful property that closed bounded sets are compact — every sequence in such a set has cluster points.
We make the set of sequences $(x_{\alpha}:\mathscr{X}_{\alpha})$ into a metric space. This is the product space $\prod_{\alpha}\mathscr{X}_{\alpha}$ with metric
$d({x},{y}){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sum_{\alpha}2^{-\alpha}\max(\|x_{\alpha}-y_{\alpha}\|,1).$ (9.1)
A sequence $({x}_{n})$ in this space is Cauchy, i.e., $\lim_{n,m\to\infty}d({x}_{n},{x}_{m})\to 0$, if and only if the equence $(x_{n,\alpha})$ of components is Cauchy for each $\alpha$.
Lemma 9.1.
  1. (a).
    $(\prod\mathscr{X}_{\alpha},d)$ is a complete metric space.
  2. (b).
    The set of projective elements is closed.
  3. (c).
    For any sequence $b_{\alpha}$, the set $\left\{{x}\in\prod\mathscr{X}_{\alpha}\,\colon\,\|x_{\alpha}\|\leq b_{\alpha}\right\}$ is compact.
Proof.
(a). According to the description just before the Lemma statement, $(x_{\alpha})$ is Cauchy if and only if each sequence $(x_{n,\alpha})_{n}$ of components is so. But, this sequence converges in $\mathscr{X}_{\alpha}$ because the latter is complete, being a finite-dimensional vector space.
(b). Assume $f_{n}\to g$ and each $f_{n}$ is projective. This means $f_{n,\alpha}\xrightarrow{n\to\infty}g_{\alpha}$ for each $\alpha$, but $f_{m,\beta}=\mathscr{P}_{\beta}f_{m,\alpha}$ for $\beta\leq\alpha$, and the RHS of this converges to $\mathscr{P}_{\beta}g_{\beta}$ by continuity of $\mathscr{P}_{\beta}$.
(c). This is a straightforward application of Tychonoff’s theorem[21, 20, 19]. ∎
For a prefiltration, $\mathscr{X}$ can be identified with a subset of the projective sequences in $\prod_{\alpha}\mathscr{X}_{\alpha}$, namely by identifying $x$ with the sequence $(\mathscr{P}_{\alpha}x)$. In this way, the metric $d$ is tranferred to $\mathscr{X}$. Moreover, if $(\mathscr{X},\mathscr{P})$ is a filtration, then $\mathscr{X}$ corresponds exactly to the projective sequences, hence becomes a complete metric space with the $d$-metric. Since the projections are bounded, Lemma 9.1(c) implies that the closed ball of norm-radius $r$ in $\mathscr{X}$, $r\cdot\overline{\mathrm{ball}}(\mathscr{X}){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\left\{x\in\mathscr{X}\,\colon\,\|x\|\leq r\right\}$, is precompact, that is, any sequence in it has a $d$-convergent subsequence. It does not generally follow that the limit is also in $r\cdot\overline{\mathrm{ball}}(\mathscr{X})$. That is true, though, for a norm filtration or SOT filtration when all projections have norm no greater than 1. Hence, in these cases, $r\cdot\overline{\mathrm{ball}}(\mathscr{X})$ is $d$-compact.
Proof.
Let $(x_{n})$ be a $d$-convergent sequence in $r\cdot\overline{\mathrm{ball}}(\mathscr{X})$. Since $\|\mathscr{P}_{\alpha}\|\leq 1$, $\|x_{n,\alpha}\|\leq r$. Since $\mathscr{X}_{\alpha}$ is finite dimensional, $x_{\alpha}=\lim_{n\to\infty}x_{n,\alpha}$ has norm not exceeding $r$. The sequence $(x_{n})$, understood as a sequence in $\mathscr{X}$, converges to $x$, either in norm of SOT, according to the case. In the first case, $\|x\|=\lim\|x_{n}\|$, and in the latter, $\|x\|\leq\lim\|x_{n}\|$ (Lemma 7.1(b)). ∎
9.1 strength of the $d$-topology
The $d$-topology is very weak. For instance, for $T\in\mathfrak{B}(\mathcal{H},\mathcal{K})$, $\mathscr{P}_{\alpha}T\mathscr{P}_{\alpha}$ amounts to a finite number of coordinates (linear functionals) of $Tv$ for a finite number of $v$ (basis of $\mathcal{H}_{\alpha}$). Thus, it is weaker than WOT. To put it another way, there are much fewer $d$-closed (or open) sets that WOT-closed, in some vague sense. Thus, it is gratifying to have the following criterion, which will be used in §10.2.
Lemma 9.2.
Let $(\mathscr{X},\mathscr{P})$ be a $\tau$-filtration ($\tau=$ norm or SOT). A $\tau$-closed subset $S$ of $\mathscr{X}$ which is mapped into itself by every projection $\mathscr{P}_{\alpha}$ is $d$-closed.
Proof.
Let $(x_{n})\subset S$ be $d$-convergent to $x$. Then, $\mathscr{P}_{\alpha}x_{n}\to\mathscr{P}_{\alpha}x$, and each $\mathscr{P}_{\alpha}x_{n}\in S$, so $\mathscr{P}_{\alpha}x\in S$. But, $\mathscr{P}_{\alpha}x\xrightarrow{\tau}x$, so $x\in S$. ∎
10 CP maps in separable setting
The main results of this section are
  1. 1.
    $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed.
  2. 2.
    Extreme CP maps and $\bm{\theta}$}-maps are the same subclass of $\mathsf{CP}(\mathcal{H},\mathcal{K})$.
  3. 3.
    Any CP operator can be (Kraus) decomposed into a countable sum of $\bm{\theta}$}-maps.
Item 1 means that if a sequence of CP maps converges in $d$-topology (or WOT, SOT, or norm), then the limit is also CP. The equivalence of item 2 above has its own interest, but is also crucial to item 3. The standard proof that Kraus decompositions exists rests on Stinespring’s theorem and a knowledge of the representations of the $C^{*}$-algebra $\mathfrak{B}(\mathcal{H})$. An explanation of how to use these results to show existence of Kraus decompositions can be found in Theorem 3.3 of [11] or Chapter 3 of [12]. Expositions of Stinespring’s theorem itself can be found in 34.7 of Ref. [23], Thm. IX.4.2 of Ref. [24], Thm. IV.3.6 of Ref. [25], or §9.2 of Ref. [26].
In comparison to the traditional method, which feels like a deus ex machina, the approach taken here is very pedestrian. We show that for any nonzero CP map $\Lambda$, we can subtract a $\bm{\theta}$}-map and have a CP map left over. Then we show that if one just keeps doing that (with some good choices), everything will converge as desired. It is worth pointing out that, even at the level of the result, existence of Kraus decompositions in the form of a sum is not obviously to be expected based on the finite-dimensional situation. From a convex analysis perspective, or considering the phenomenon of continuous spectrum, some sort of integral would not be surprising.
The treatment of the GKSL problem in §11 relies on results of §10.2. Otherwise, it is independent of this section.
10.1 Partial orders on $\mathsf{HP}(\mathcal{H})$
How do we make sense of a putative Kraus decomposition $\sum_{i=1}^{\infty}{\bm{\theta}}(A_{i})$ with infinitely many terms? As some sort of monotonic approach of the partial sums to the limit. We need to carefully investigate the order corresponding to “monotonic”, but it seems advisable to start with $\mathcal{B}^{1}(\mathcal{H})$. The partial order $\leq$ on this space was defined way back in §3.1.2. $\rho\leq\sigma$ means that $\sigma-\rho$ is in the cone $\mathrm{Pos}(\mathcal{H})$, and $\rho<\sigma$ means that, in addition, $\rho\neq\sigma$. Similarly, the cones $\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathsf{Mon}(\mathcal{H},\mathcal{K})$ induce partial orders on $\mathsf{HP}(\mathcal{H},\mathcal{K})$. $\Lambda\leq\Gamma$ means that $\Gamma-\Lambda$ is monotone. The corresponding strict partial order $\Lambda<\Gamma$ means that, in addition, $\Lambda\neq\Gamma$. We may, as convenient, write $\geq$ and >> for the reverse order. So, $\Lambda\geq 0$ is synonymous with $\Lambda\in\mathsf{Mon}(\mathcal{H},\mathcal{K})$ except that it isn’t explicit about $\mathcal{H}$ and $\mathcal{K}$. This order can be unpacked in terms of that on $\mathcal{B}^{1}(\mathcal{H})$. By definition of $\mathsf{Mon}$, $\Lambda\leq\Gamma$ means that $\Lambda\rho\leq\Gamma\rho$ for every $\rho\colon\mathrm{Pos}(\mathcal{H})$. The two occurences of $\leq$” refer to partial orders on different spaces. $\Lambda\precsim\Gamma$ means that $\Gamma-\Lambda$ is CP, with other variations as for $\leq$. Since $\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})$,
$\Lambda\precsim\Gamma\quad\Rightarrow\quad\Lambda\leq\Gamma.$ (10.1)
[…]
Lemma 10.1.
$0\leq\Lambda\leq\Gamma$ implies that $\|\Lambda\|\leq\|\Gamma\|$.
Proof.
Choose $\rho\geq 0$ with $\|\rho\|_{1}=\operatorname{\mathrm{Tr}}\rho=1$ such that $\|\Lambda\rho\|_{1}=\operatorname{\mathrm{Tr}}\Lambda\rho>\|\Lambda\|-\epsilon$. Then, $\|\Gamma\|\geq\operatorname{\mathrm{Tr}}\Gamma\rho\geq\operatorname{\mathrm{Tr}}\Lambda\rho$. ∎
10.2 $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed
Lemma 10.2.
$\mathsf{Mon}(\mathcal{H},\mathcal{K})$ is WOT-closed.
Proof.
Suppose $\Phi$ is not monotone. Then there is a pair $(B,\rho)\in\mathfrak{B}(\mathcal{K})^{+}\times\mathcal{B}^{1}(\mathcal{H})^{+}$ witnessing that fact via $\langle B,\Phi\rho\rangle\not\in[0,\infty)$, so $\{\Gamma:\langle B,\Gamma\rho\rangle\not\in[0,\infty)\}$ is a WOT neighborhood of $\Phi$ free of monotone superoperators. ∎
Lemma 10.3.
$\Gamma\mapsto\Gamma\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}$ is WOT/WOT continuous.
Proof.
Use the realization of $\rho\colon\mathcal{B}^{1}(\mathcal{K}\otimes{\mathbb{C}}^{N})$ as an $N\times N$ matrix with entries in $\mathcal{B}^{1}(\mathcal{K})$. Then $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}$ operates on it by application of $\Lambda$ to each entry. $\langle{B}\mid{\Gamma_{n}\rho}\rangle\to\langle{B}\mid{\Lambda\rho}\rangle$ is equivalent to $\langle{B_{ij}}\mid{\Gamma_{n}\phi_{ij}}\rangle\to\langle{B_{ij}}\mid{\Lambda\phi_{ij}}\rangle$ for each $i,j$. But this latter is implied by $\Gamma_{n}\stackrel{{\scriptstyle WOT}}{{\longrightarrow}}\Lambda$ since there are only finitely many $i,j$. ∎
Lemma 10.4.
$\mathsf{CP}(\mathcal{H},\mathcal{K})$ is WOT-closed.
Proof.
If $(\Lambda_{n})$ is a net in $\mathsf{CP}(\mathcal{H},\mathcal{K})$, WOT-convergent to $\Lambda$, then $\Lambda_{n}\otimes\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}$ in $\mathsf{Mon}(\mathcal{H},\mathcal{K})$ is WOT-convergent to $\Lambda\otimes\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}$ (Lemma 10.3), which is monotone by Lemma 10.2. ∎
Proposition 10.5.
$\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed.
Proof.
$\mathsf{CP}(\mathcal{H},\mathcal{K})$ is WOT closed by Lemma 10.4. It is also stable under projection: $\mathscr{P}_{\alpha}(\Lambda)$ is CP if $\Lambda$ is so. The conclusion then follows by Lemma 9.2. ∎
Proposition 10.6.
[…]
All this means is that if $\Lambda_{n}\precsim\Gamma_{n}$, $\Lambda_{n}\xrightarrow{d}\Lambda$, and $\Gamma_{n}\xrightarrow{d}\Gamma$, then $\Lambda\precsim\Gamma$. It follows immediately from Prop. 10.5 because what those relations mean is $\Gamma_{n}-\Lambda_{n}\in\mathsf{CP}(\mathcal{H},\mathcal{K})$. Conversely, Prop. 10.5 is also an immediate corollary of Prop. 10.6.
10.3 The monoidal functor $\bm{\theta}$
§4.2 showed that — for finite-dimensional Hilbert spaces — $\mathsf{Hilb}$ and $\mathsf{CP}$ are monoidal categories under tensor product and $\bm{\theta}$} is a monoidal functor $\mathsf{Hilb}\rightarrow\mathsf{CP}$. There is a potential difficulty in extending this to separable Hilbert spaces. According to Def. 3.4, $\Lambda$ is CP if $\Lambda\otimes\operatorname{\mathbbm{1}}_{{\mathbb{C}}^{N}}$ is monotone for any natural number $N$. In the finite-dimensional setting, this was enough to guarantee (Prop. 3.1) that $\Lambda\otimes\Gamma$ is CP as soon as $\Lambda$ and $\Gamma$ are so, because they operate on finite-dimensional spaces. Is it necessary to strengthen the defintion of CP in the separable setting? In fact no, as §10.3.1 shows. §10.3.2 establishes some useful properties of $\bm{\theta}$}.
10.3.1 Tensor products and the CP concept
First, we must identify the space on which $\Lambda\otimes\Gamma$ operates, assuming $\Lambda\colon\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$ and $\Gamma\colon\mathfrak{B}(\mathcal{B}^{1}(\mathcal{K}))$. In contrast to Hilbert spaces, the choice of norm on an algebraic tensor product of Banach spaces, hence its completion, is often not canonical[27, 28]. Fortunately, in our case, we can identify $\mathcal{B}^{1}(\mathcal{H})\otimes\mathcal{B}^{1}(\mathcal{K})$ (more precisely, its completion) with $\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{K})$, the tensor product in this last expression being the ordinary Hilbert one.
Lemma 10.7.
If $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathcal{N}$ is separable, then $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})}$ is monotone.
Proof.
$\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{K}\otimes\mathcal{N}$ are Hilbert spaces, so the filtration theory applies to $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{N}),\mathcal{B}^{1}(\mathcal{K}\otimes\mathcal{N}))$ as usual. However, we choose a factorized form of the basic filtrations on $\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{K}\otimes\mathcal{N}$, e.g., $P_{\alpha}^{\mathcal{H}\otimes\mathcal{K}}=P_{\alpha}^{\mathcal{H}}\otimes P_{\alpha}^{\mathcal{K}}$. Now, because $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{K}))$ is a SOT filtration (Ex. 8.5), $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})}$ is the strong limit of $\mathscr{P}_{\alpha}(\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})})=\mathscr{P}_{\alpha}(\Lambda)\otimes\mathscr{P}_{\alpha}(\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})})$, which can be identified with
$\mathscr{P}_{\alpha}(\Lambda)\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N}_{\alpha})}.$ (10.2)
The important point is that $\mathcal{N}_{\alpha}$ is finite-dimensional. Since $\mathscr{P}_{\alpha}(\Lambda)=\theta(P_{\alpha})\circ\Lambda$ is CP, (10.2) is monotone. By Lemma 10.4, so is $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})}$. ∎
10.3.2 Some properties of $\bm{\theta}$
We collect here a few properties of $\bm{\theta}$} that will be needed in this section.
Lemma 10.8.
  1. (a).
    $\|{\bm{\theta}}(A)\|=\|A\|^{2}$,
  2. (b).
    $\mathscr{P}_{\alpha}\circ{\bm{\theta}}={\bm{\theta}}\circ\mathscr{P}_{\alpha}$,
  3. (c).
    $\bm{\theta}$} is norm/norm and dd/$d$ continuous, and SOT/SOT continuous on bounded sets.
Proof.
(a). On the one hand, $\|A\rho A^{\dagger}\|_{1}\leq\|A\|\|\rho A|_{1}\leq\|A\|^{2}\|\rho|_{1}$. On the other, let $h$ be a unit vector nearly saturating the norm of $A$: $\|Ah\|\simeq\|A\|$. Then $\|A{\bm{\pi}}(h)A^{\dagger}\|_{1}=\|{\bm{\pi}}(Ah)\|_{1}=\|Ah\|^{2}\simeq\|A\|^{2}$.
(b). $\mathscr{P}_{\alpha}({\bm{\theta}}(A))\cdot\rho=\mathscr{P}_{\alpha}({\bm{\theta}}(A)(\mathscr{P}_{\alpha}\rho))=P_{\alpha}(AP_{\alpha}\rho P_{\alpha}A^{\dagger})P_{\alpha}=(P_{\alpha}AP_{\alpha})\rho(P_{\alpha}A^{\dagger}P_{\alpha})={\bm{\theta}}(P_{\alpha}AP_{\alpha})\cdot\rho={\bm{\theta}}(\mathscr{P}_{\alpha}(A))\cdot\rho.$
(c). $d{\bm{\theta}}(A)\cdot B=B\,\square\,A^{\dagger}+A\,\square\,B^{\dagger}$ is a bounded $\mathbb{R}$}-linear operator. So, $\bm{\theta}$} is not only norm-continuous, but norm-differentiable. By part (b), $\mathscr{P}_{\alpha}({\bm{\theta}}(A))={\bm{\theta}}(\mathscr{P}_{\alpha}(A))$. Thus, by the just-proven norm continuity, dd/$d$ continuity follows. Finally, if $\rho=\sum_{i=1}^{\infty}{\bm{\pi}}(h_{i})$, then for $B\in\overline{\mathrm{ball}}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$, $B\rho B^{\dagger}\approx A\rho A^{\dagger}$ as soon as $Bh_{i}\approx Ah_{i}$ for $i=1,\ldots,N$ for some $N$. This shows SOT/SOT continuity on bounded sets. ∎
10.4 Extreme CP map = $\bm{\theta}$-map
Corollary 10.10 says that $\bm{\theta}$}-maps are extreme, and Lemma 10.11 immediately implies that a nonzero extreme $\Lambda$ is a $\bm{\theta}$}-map. [If $0\prec{\bm{\theta}}(A)\precsim\Lambda$ and $\Lambda$ is extreme, then necessarily $\Lambda={\bm{\theta}}(A)$.] Note, however, that the Lemma is stronger than that; the extra strength will be needed in §10.6.
Lemma 10.9.
If $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ is not an extreme CP map, then for sufficiently large $\alpha$, $\mathscr{P}_{\alpha}\Lambda$ is not extreme.
Proof.
Suppose $\Lambda=\Gamma+\Gamma^{\prime}$ is a non-trivial CP decomposition. Then, $\mathscr{P}_{\alpha}\Lambda=\mathscr{P}_{\alpha}\Gamma+\mathscr{P}_{\alpha}\Gamma^{\prime}$; the terms on the right-hand side tend to $\Gamma$ and $\Gamma^{\prime}$, respectively, in SOT. Now, suppose for a contradiction that along some subsequence, $\mathscr{P}_{\alpha}\Gamma=c_{\alpha}\mathscr{P}_{\alpha}\Gamma^{\prime}$ with $0<c_{\alpha}<\infty$. Then by SOT convergence, it is not possible that the $c_{\alpha}$ are unbounded. Therefore, there is a subsequence along which $c_{\alpha}$ converges to $c\geq 0$. But then, $\Gamma=c\Gamma^{\prime}$. ∎
Corollary 10.10.
$\bm{\theta}$}-maps are extreme CP maps.
Proof.
Immediate from $\mathscr{P}_{\alpha}({\bm{\theta}}(A))={\bm{\theta}}(\mathscr{P}_{\alpha}(A))$, the finite-dimensional theory, and Lemma 10.9. ∎
Lemma 10.11.
[…]
Proof.
Given $0\prec\Lambda$, there is $h$ such that $\Lambda{\bm{\pi}}(h)$ has a nontrivial orthogonal decomposition ${\bm{\pi}}(k)+\cdots$. Take some filtration with $\mathcal{H}_{1}=\operatorname{\mathrm{span}}\{h,k\}$. Then, $(\mathscr{P}_{\alpha}\Lambda){\bm{\pi}}(h)=\mathscr{P}_{\alpha}(\Lambda{\bm{\pi}}(h))={\bm{\pi}}(k)+\cdots$. By Lemma 10.12 below, Kraus decompositions $\mathscr{P}_{\alpha}\Lambda=\sum_{i}{\bm{\theta}}(A^{\alpha}_{i})$ of the finite-dimensional projections of $\Lambda$ may be chosen so that $\left\langle{k}\middle|{A_{1}^{\alpha}h}\right\rangle=\|k\|$. Since $\Lambda\succsim{\bm{\theta}}(A_{1}^{\alpha})$, the sequence $(A_{1}^{\alpha})$ is bounded and therefore has a $d$ cluster point, $A$, which is nonzero because $\left\langle{k}\middle|{Ah}\right\rangle=\|k\|$. Summing up, $\Lambda\succsim{\bm{\theta}}(A)\succ 0$. ∎
Lemma 10.12.
Let $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ where $\mathcal{H}$, $\mathcal{K}$ are finite-dimensional, and let $\Lambda{\bm{\pi}}(h)$ have orthogonal decomposition $\Lambda{\bm{\pi}}(h)={\bm{\pi}}(k)+\cdots$. Then there exists a choice of $A_{i}$ in the Kraus decomposition $\Lambda=\sum_{i}{\bm{\theta}}(A_{i})$ such that
$\left\langle{k}\middle|{A_{i}h}\right\rangle=\begin{cases}\|k\|,&i=1\\ 0,&\text{otherwise}.\end{cases}$ (10.3)
Proof.
For convenience, scale $h$ so that $\|k\|=1$. Let $\Lambda=\sum_{j}{\bm{\theta}}(B_{j})$ be any Kraus decomposition. Then
$\sum_{j}\left|\left\langle{k}\middle|{B_{j}h}\right\rangle\right|^{2}=\sum_{j}\left\langle{{\bm{\pi}}(k)}\middle|{{\bm{\theta}}(B_{j}){\bm{\pi}}(h)}\right\rangle=\left\langle{{\bm{\pi}}(k)}\middle|{\Lambda{\bm{\pi}}(h)}\right\rangle=1$ (10.4)
To change the decomposition basis, we define $A_{i}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sum U_{ij}B_{j}$, with $U$ a unitary matrix. It may be chosen with first row $U_{1j}=\left\langle{B_{j}h}\middle|{k}\right\rangle$, since by (10.4), $\sum_{j}|U_{1j}|^{2}=1$. By the same token, $\left\langle{k}\middle|{A_{1}h}\right\rangle=\sum_{j}U_{1j}\left\langle{k}\middle|{B_{j}h}\right\rangle=1$. Finally, however the rest of $U$ is filled out, (10.4) with $B$’s replaced by $A$’s shows that $\left\langle{k}\middle|{A_{i}h}\right\rangle=0$ for $i\neq 1$. ∎
10.5 Order closedness of $\mathsf{CP}(\mathcal{H},\mathcal{K})$
Recall that a $\leq$-directed set $D$ is a set such that, for any $x,y\in D$, there is $z\in D$ such that $x\leq z$ and $y\leq z$. A chain, which is a totally ordered set (everything is comparable) is a special case.
Lemma 10.13.
A $\leq$-directed set $D$ in $\mathcal{B}^{1}(\mathcal{H})$ which is bounded above has a supremum and converges to it in norm.
Proof.
Without loss, assume $D\subset\mathrm{Pos}(\mathcal{H})$, otherwise translate $D$ to move some member to zero, so that the part of $D$ above it becomes positive.
By boundedness, the functions $\left\{x\mapsto\left\langle{x}\middle|{\rho x}\right\rangle\,\colon\,\rho\in D\right\}$ certainly have an upper bound $f(x)$. For any $x$ and $\epsilon>0$, there is a member of $D$ satisfying the predicate $Q_{\epsilon,x}(\rho)\equiv(\left\langle{x}\middle|{\rho x}\right\rangle>f(x)-\epsilon$). The significance of directedness is that any given finite number of such predicates is satisfied by some member of $D$. This allows one to show that a positive trace-class linear operator $\sigma$ can be recovered from $f$ by polarization: $\left\langle{x}\middle|{\sigma y}\right\rangle{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\tfrac{1}{4}\sum_{k}(-i)^{k}f(x+i^{k}y)$. E.g., we can approximate all four terms $f(x+x^{\prime}+i^{k}y)$ by a member of $D$ to show conjugate linearity in $x$. So, $\sigma$ has an orthogonal decomposition $\sigma=\sum c_{n}{\bm{\pi}}({h}_{n})$, with unit vectors ${h}_{n}$, and $\sum c_{n}<\infty$. Since any $\rho\in D$ is bounded by $\sigma$, $\|\sigma-\rho\|=\operatorname{\mathrm{Tr}}(\sigma-\rho)$, so a trace-norm approximation is achieved by satisfying some finite number of $Q_{\epsilon,h_{n}}$’s. ∎
Lemma 10.14.
Let $D$ be a $\leq$-directed set in $\mathsf{HP}(\mathcal{H},\mathcal{K})$, bounded above. Then $\sup D$ exists and $D$ converges to it in SOT. An analogous statement holds for a $\geq$-directed set.
Proof.
The overall strategy of the proof is much like that of Lemma 10.13. Given $\rho\colon\mathcal{B}^{1}(\mathcal{H})$, $D(\rho){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\{\Gamma\rho\,:\,\Gamma\in D\}$ is a directed set in $\mathcal{B}^{1}(\mathcal{H})$, bounded above, exactly the subject of Lemma 10.13. Define a function $f(\rho){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sup(D(\rho))$; $D(\rho)$ converges to $\sup(D(\rho))$ in norm, so $D$ converges to $f$ in SOT. It only remains to used directedness to show that $f$ is linear: Given $\rho,\rho^{\prime}$, there is $\Lambda\in D$ such that $f(\rho+c\sigma)\approx\Lambda(\rho+c\sigma)=\Lambda\rho+c\Lambda\sigma\approx f(\rho)+cf(\sigma)$. ∎
Proposition 10.15.
If $D$ is a bounded-above $\leq$-directed set contained in $\mathsf{CP}(\mathcal{H},\mathcal{K})$, then so is its supremum. An analogous statement holds for a $\geq$-directed set.
Proof.
By Lemma 10.14, $D$ has a supremum in $\mathsf{HP}(\mathcal{H},\mathcal{K})$ and converges to it in SOT, but $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is SOT-closed by Lemma 10.4. ∎
10.6 Every CP map has Kraus decompositions
Lemma 10.11 is the key to Kraus decomposition in our approach. Given $0\prec\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$, we can subtract some ${\bm{\theta}}(A_{1})$ and still have a CP operator. Then subtract ${\bm{\theta}}(A_{2})$, and so forth until there is nothing left. All the terms we subtracted are a Kraus decomposition.
Let us set up some auxiliary machinery to do that more carefully. Referring to some filtration on $\mathcal{H}$, define $\rho_{0}=c_{1}P_{1}+c_{2}(P_{2}-P_{1})+c_{3}(P_{3}-P_{2})+\cdots$, with $c_{1}>c_{2}>\cdots>0$ chosen so that $\operatorname{\mathrm{Tr}}\rho_{0}<\infty$. This can be done, since each $P_{n+1}-P_{n}$ is a finite-dimensional orthoprojector. Next, define a linear functional $f\colon\mathsf{HP}(\mathcal{H},\mathcal{K})\rightarrow{\mathbb{R}}$ by $f(\Gamma)=\operatorname{\mathrm{Tr}}\Gamma\rho_{0}$. Clearly, $0\leq\Gamma$ implies $0\leq f(\Gamma)$. Furthermore, restricted to $\mathcal{H}_{n}$, $\rho_{0}\geq c_{n}\operatorname{\mathbbm{1}}_{\mathcal{H}_{n}}>0$, so that for such $\Gamma$, $f(\Gamma)=0$ implies $\mathscr{P}_{n}\Gamma=0$ for all $n$ and therefore $\Gamma=0$. So, $\Lambda\leq\Gamma$ implies $f(\Lambda)\leq f(\Gamma)$ and $\Lambda<\Gamma$ implies $f(\Lambda)<f(\Gamma)$.
Now, suppose we are given $0\prec\Lambda{\kern 1.99997pt}={\kern-5.0pt}:{\kern 1.99997pt}\Lambda_{0}$. We are going to define inductively a decreasing sequences $(\Lambda_{n}\colon\mathsf{CP}(\mathcal{H},\mathcal{K}))$ and a sequence $(A_{n}\colon\mathfrak{B}(\mathcal{H},\mathcal{K}))$. Given $\Lambda_{n}$, let $S_{n+1}=\{B\,:\,{\bm{\theta}}(B)\precsim\Lambda_{n}\}$, and choose $A_{n+1}$ in $S_{n+1}$ such that
$f({\bm{\theta}}(A_{n+1}))>\sup\{f({\bm{\theta}}(B))\,:\,B\in S_{n+1}\}-\tfrac{1}{n+1}.$ (10.5)
Then, define $\Lambda_{n+1}=\Lambda_{n}-{\bm{\theta}}(A_{n+1})$. For convenience, also define $K_{0}=0$, $K_{n+1}=K_{n}+{\bm{\theta}}(A_{n+1})$.
Now, we have $\Lambda=\Lambda_{n}+K_{n}$, so $f(\Lambda)=f(\Lambda_{n})+f(K_{n})$. In particular, $f(K_{n})$ is a sum of positive numbers not exceeding $f(\Lambda)$. Therefore $f({\bm{\theta}}(A_{n}))\to 0$ and
$\sup\{f({\bm{\theta}}(B))\,:\,B\in S_{n}\}\to 0,\text{ as }n\to\infty.$ (10.6)
Prop. 10.15 shows that $\Lambda_{\infty}=\lim\Lambda_{n}$ and $K_{\infty}=\lim K_{n}$ exist in $\mathsf{CP}(\mathcal{H},\mathcal{K})$. Suppose, for a contradiction, that $0\prec\Lambda_{\infty}$. Then, by Lemma 10.11, as usual, there is some $0\prec{\bm{\theta}}(A)\precsim\Lambda_{\infty}$, and $f({\bm{\theta}}(A))>0$. But $A$ belongs to every $S_{n}$, and this is impossible by (10.6). Therefore, $K_{\infty}=\Lambda$ and the $A_{n}$ are the elements of a Kraus decomposition.
11 GKSL problem in separable setting
Recall from §5 the map
$\begin{array}[]{ccl}\mathsf{Lind\,{\mathcal{H}}}&\xrightarrow{L}&\mathsf{HP}({\mathcal{H}})\\ (\Psi,G,H)&\mapsto&\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;].\end{array}$ (11.1)
§5.4 showed that $L(\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}})$ is a cone in $\mathsf{cp}^{+}(\mathcal{H})$, which was called the $L$-cone. §5.4 also showed that a continuously varying $\Lambda(t)$ in the $L$-cone generates a Markovian evolution via the equation $\dot{\mathcal{E}}(t)=\Lambda(t)\mathcal{E}(t)$. The proofs of these assertions given there work for the separable case, with only small emendations involving the operator norm. In the definition of forward-tangent (or tangent), it is necessary to specify that “o(t)o(t)” means in operator norm. And, we need norm continuity of $\Lambda(t)$.
The second part of the project of §5 showed that the $L$-cone is actually $\mathsf{cp}^{+}(\mathcal{H})$, and that Lindblad parametrizers exist. The proof of this converse direction relied on the Jamiołkowski transform, hence we cannot transfer it directly to the separable setting. The body of this section is devoted to establishing the existence of such parametrizers. Recall that a Lindblad parametrizer (Def. 5.3) is a bounded linear map $\Delta\colon\mathsf{HP}(\mathcal{H})\rightarrow\mathsf{Lind\,{\mathcal{H}}}$ which is right-inverse to $L$, i.e., $L(\Delta\Lambda)=\Lambda$ and such that if $\Lambda$ is CP, then the first component of $\Delta\Lambda$ is CP.
§11.1 establishes a technical result for finite-dimensional Hilbert spaces which plays a crucial rôle. §11.2 gives some other technical results. Finally, §11.3 shows that the sequence of Lindblad parametrizers over the $\mathcal{H}_{\alpha}$ has cluster points which are Lindblad parametrizers over $\mathcal{H}$.
11.1 Dimension-independent bound on the minimal Lindblad parametrizer
Lemma 11.1.
The operator norms of both $L$ and $\Delta^{\scriptstyle{min}}$ for finite-dimensional Hilbert space are bounded independently of dimension.
The proof method borrows a trick from Thm. 5 of Lindblad[2]. We consider functions $f(U)$ of a unitary variable $U$, and average over the unitary group $\mathsf{U}(\mathcal{H})$, an operation denoted $\langle f(U)\rangle$. In other words $\langle\cdot\rangle$ is integration with respect to a measure on $\mathsf{U}(\mathcal{H})$ which is normalized ($\langle A\rangle=A$ if $A$ does not depend on $U$) and invariant ($\langle f(VU)\rangle=\langle f(U)\rangle$ for fixed unitary $V$). An explicit construction will not be needed.
Consider, in particular, a function of form ${\bm{\theta}}(U)A$ for $A\colon\mathfrak{B}(\mathcal{H})$. $\langle{\bm{\theta}}(U)A\rangle$ commutes with all unitaries, since $V\langle{\bm{\theta}}(U)A\rangle V^{\dagger}=\langle{\bm{\theta}}(VU)A\rangle$, and therefore $\langle{\bm{\theta}}(U)\rangle$ is proportional to orthoprojection onto $\operatorname{\mathbbm{1}}_{\mathcal{H}}$ [a form of Schur’s Lemma]. Since ${\bm{\theta}}(U)\operatorname{\mathbbm{1}}_{\mathcal{H}}=\operatorname{\mathbbm{1}}_{\mathcal{H}}$, the proportionality constant is one:
$\langle{\bm{\theta}}(U)\rangle={\operatorname{\mathbbm{1}}_{\mathcal{H}}}\tfrac{\operatorname{\mathrm{Tr}}}{\dim\mathcal{H}}=\hat{\bm{\pi}}[\operatorname{\mathbbm{1}}_{\mathcal{H}}].$ (11.2)
Proof of Lemma 11.1.
Take $\Lambda:\mathsf{HP}(\mathcal{H})$, with $\Delta^{\scriptstyle{min}}(\Lambda)=(\Psi,G,H)$.
$\begin{split}\langle\Lambda(U)\cdot U^{-1}\rangle=&\langle\Psi(U)\cdot U^{-1}\rangle-\langle[G,U]_{+}\cdot U^{-1}\rangle-\langle[iH,U]\cdot U^{-1}\rangle\\ =&{\bm{\{}}{\operatorname{\mathrm{Tr}}A=0\Rightarrow\langle{\bm{\theta}}(A)U\cdot U^{-1}\rangle=A\langle UA^{\dagger}U^{-1}\rangle=0}{\bm{\}}}\\ &-\langle[G,U]_{+}\cdot U^{-1}\rangle-\langle[iH,U]\cdot U^{-1}\rangle\\ =&-\left(G+\tfrac{\operatorname{\mathrm{Tr}}G}{\dim\mathcal{H}}\operatorname{\mathbbm{1}}_{\mathcal{H}}\right)-iH.\end{split}$ (11.3)
$\|\langle\Lambda(U)\cdot U^{-1}\rangle\|\leq\|\Lambda\|$ because we’re averaging over $U$ and $\|U\|=1$. Since the norm of any operator is at least as large as that of its hermitian or antihermitian part,
$\|G+\tfrac{\operatorname{\mathrm{Tr}}G}{\dim\mathcal{H}}\operatorname{\mathbbm{1}}_{\mathcal{H}}\|,\,\|H\|\leq\|\Lambda\|.$ (11.4)
Now, assume $\|G\|$ is an eigenvalue of $G$ (otherwise, $\|G\|$ is one, and we work with G-G), and abbreviate $c=\tfrac{1}{\dim\mathcal{H}}\operatorname{\mathrm{Tr}}G$. If $c>0$, then clearly $\|G\|\leq\|G+c\operatorname{\mathbbm{1}}_{\mathcal{H}}\|$. If $c\leq 0$, then $G$ must have a non-positive eigenvalue, so the largest eigenvalue of $G-|c|\operatorname{\mathbbm{1}}_{\mathcal{H}}$ is $\|G\|-|c|$ and the smallest is no greater than |c|-|c|. No matter the value of $c$, one of these is at least $\tfrac{\|G\|}{2}$ in absolute value. Therefore, $\|G\|\leq 2\|G+c\operatorname{\mathbbm{1}}_{\mathcal{H}}\|\leq 2\|\Lambda\|$. Finally,
$\|\Psi\|\leq\|\Lambda\|+\|[G,\;\cdot\;]_{+}\|+\|[iH,\;\cdot\;]\|\leq\|\Lambda\|+4\|\Lambda\|+2\|\Lambda\|.$ (11.5)
11.2 Projector juggling
The operator $\Delta^{\scriptstyle{min}}_{\alpha}$ a priori defined on $\mathsf{HP}(\mathcal{H}_{\alpha})$ with range in $\mathsf{Lind\,{\mathcal{H}_{\alpha}}}$ has a natural promotion to an operator $\mathsf{HP}(\mathcal{H})\rightarrow\mathsf{Lind\,{\mathcal{H}}}$. $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is the result of projecting $\Lambda$ into $\mathsf{HP}(\mathcal{H}_{\alpha})$, applying $\Delta^{\scriptstyle{min}}_{\alpha}$ as originally defined, and then regarding the result as being in $\mathsf{Lind\,{(}}\mathcal{H})$ instead. There is no real ambiguity in denoting this projected operator by the same symbol. So, we have
$\Delta^{\scriptstyle{min}}_{\alpha}=\mathscr{P}_{\alpha}\circ\Delta^{\scriptstyle{min}}_{\alpha}\circ\mathscr{P}_{\alpha}.$ (11.6)
$L\colon\mathsf{Lind\,{\mathcal{H}}}\rightarrow\mathsf{HP}(\mathcal{H})$ is defined by $L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$, just as in finite dimensions. Since the formula is the same, $(L\circ\Delta^{\scriptstyle{min}}_{\alpha})\Gamma=\Gamma$ if $\Gamma\in\mathsf{HP}(\mathcal{H}_{\alpha})$. So, if we replace $\Gamma$ in the preceding equation by $\mathscr{P}_{\alpha}\Lambda$, we deduce
$\mathscr{P}_{\alpha}\circ L\circ\Delta^{\scriptstyle{min}}_{\alpha}=\mathscr{P}_{\alpha}.$ (11.7)
The following will also be extremely useful.
$\mathscr{P}_{\alpha}(LZ){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}$ $\mathscr{P}_{\alpha}\circ LZ\circ\mathscr{P}_{\alpha}$ (11.8a)
$=$ $\mathscr{P}_{\alpha}\circ L(\mathscr{P}_{\alpha}(Z))$ (11.8b)
$=$ $L(\mathscr{P}_{\alpha}(Z))\circ\mathscr{P}_{\alpha}$ (11.8c)
Proof.
For $Z$ of form $(\Psi,0,0)$, this is trivial since $\mathscr{P}_{\alpha}(\Psi,0,0)=(\mathscr{P}_{\alpha}\Psi,0,0)$ and $L(\Psi,G,H)=\Psi$. By linearity, it only remains to verify the displayed equations for $Z$ of form $(0,G,0)$ or $(0,0,H)$ This reduces to the following directly verifiable equations:
$\begin{split}\mathscr{P}_{n}([T,\;\cdot\;]_{\pm})=&\mathscr{P}_{n}\circ[T,\;\cdot\;]_{\pm}\circ\mathscr{P}_{n}\\ =&\mathscr{P}_{n}\circ[\mathscr{P}_{n}(T),\;\cdot\;]_{\pm}\\ =&[\mathscr{P}_{n}(T),\;\cdot\;]_{\pm}\circ\mathscr{P}_{n}.\end{split}$ (11.9)
11.3 Lindblad parametrizations and parametrizers exist
The main idea now is to show existence of Lindblad parametrizations and parametrizers by using the sequence $(\Delta^{\scriptstyle{min}}_{\alpha})_{\alpha}$ as some sort of approximation to a Lindblad parametrizer. For it to fulfill that role, we need the properties established in the first couple of subsections.
11.3.1 The first component of $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP when $\Lambda$ is $\mathsf{cp}^{+}$
Lemma 11.2.
If $\Lambda$ is $\mathsf{cp}^{+}$, then the first component of $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP.
Proof.
Because $\mathscr{P}_{\alpha}$ is a contraction, if $\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}}$, then $\mathscr{P}_{\alpha}\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H}_{\alpha})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}_{\alpha}}$. Therefore, as an element of $HP(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})$, $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ has CP first component. But it is still CP regarded as an element of $\mathsf{HP}(\mathcal{H})$. ∎
11.3.2 The sequence $(\Delta^{\scriptstyle{min}}_{\alpha})$ is an approximate right-inverse of $L$
Lemma 11.3.
$L\circ\Delta^{\scriptstyle{min}}_{\alpha}(\Lambda)\xrightarrow{SOT}\Lambda$ as $\alpha\to\infty$.
Proof.
We need the following calculation:
[…]
Now, since (Lemma 11.1) $L$ and $\Delta^{\scriptstyle{min}}_{\alpha}$ are bounded uniformly in $\alpha$, it suffices to show that $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)\cdot A\rightarrow\Lambda A$ for $A$ in the set $\cup_{m}\mathcal{B}^{1}(\mathcal{H}_{m})$ dense in $\mathcal{B}^{1}(\mathcal{H})$. Thus, take AA in $\mathcal{B}^{1}(\mathcal{H}_{\beta})$. As soon as $\beta\leq\alpha$, $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)A=(\mathscr{P}_{\alpha}\Lambda)A$, by (11.10). But, $\mathscr{P}_{\alpha}\Lambda\xrightarrow{SOT}\Lambda$, and the conclusion follows immediately. ∎
11.3.3 Bringing in the $d$-metrics
Eventually, we must work with the $d$-metrics because that is the only topology we have with respect to which every bounded sequence has cluster points.
Lemma 11.4.
$L\colon\mathsf{Lind\,{\mathcal{H}}}\rightarrow\mathsf{HP}(\mathcal{H})$ is dd/$d$ continuous.
Proof.
Let $Z_{n}\xrightarrow{d}Z$ in $\mathsf{Lind\,{\mathcal{H}}}$. By definition (of $d$-convergence), this means that $\mathscr{P}_{\alpha}(Z_{n})\to\mathscr{P}_{\alpha}(Z)$ for each $\alpha$. Then, applying $L$ and composing with $\mathscr{P}_{\alpha}$, $L(\mathscr{P}_{\alpha}(Z_{n}))\circ\mathscr{P}_{\alpha}\to L(\mathscr{P}_{\alpha}(Z))\circ\mathscr{P}_{\alpha}$. Lemma 11.8c shows that this is the same as $\mathscr{P}_{\alpha}(LZ_{n})\to\mathscr{P}_{\alpha}(LZ)$, which, again by definition, is just $LZ_{n}\xrightarrow{d}LZ$. Therefore $Z$ is a Lindblad parametrization.
It remains to show that the first component of $Z$ is CP if $\Lambda\in\mathsf{cp}^{+}(\mathcal{H})$. This follows immediately from $d$-closure of $\mathsf{CP}(\mathcal{H})$ (Prop. 10.5) and the fact that the first component of each $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP. ∎
11.3.4 Individual Lindblad parametrizations
Lemma 11.5.
Given $\Lambda\colon\mathsf{HP}(\mathcal{H})$, every $d$-cluster-point of $(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)$, of which there is at least one, is a Lindblad parametrization of $\Lambda$.
Proof.
$(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)$ is bounded by Lemma 11.1, hence has $d$-cluster-points (see §9). Let $Z$ be one such. By Lemma 11.4, $LZ$ is a $d$-cluster point of the sequence $(L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda))$. But Lemma 11.3 says that only $\Lambda$ can be such a cluster point.
Finally, if $\Lambda$ is $\mathsf{cp}^{+}$, then the first component of $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP, according to Lemma 11.2. Since $\mathsf{CP}(\mathcal{H})$ is $d$-closed (Prop. 10.5), every $d$-cluster point of $(\Delta^{\scriptstyle{min}}_{\alpha})$ also has this property. ∎
11.3.5 Full Lindblad parametrizer
Proposition 11.6.
The sequence $(\Delta^{\scriptstyle{min}}_{\alpha})$ has a cluster point which is a Lindblad parametrizer.
Proof.
Let $\Lambda_{1},\Lambda_{2},\ldots$ be a countable dense subset of $\mathsf{HP}(\mathcal{H})$. The rough idea here is to use Lemma 11.5, or better its proof, simultaneously for every $\Lambda_{n}$.
In $\mathsf{Lind\,{\mathcal{H}}}$ equipped with the $d$-metric, $\left\{\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}\,\colon\,\alpha\in{\mathbb{N}}\right\}$ is a precompact set. The product space $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$ can be turned into a metric space just as in §9, such that products of compact sets are compact. Thus, the sequence $\alpha\mapsto\prod_{n}\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}$ is precompact in $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$. $L^{\mathbb{N}}$ carries any cluster point of it into a cluster point of the sequence $\alpha\mapsto\prod_{n}L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n})$ in $\mathsf{HP}(\mathcal{H})^{\mathbb{N}}$, by Lemma 11.4.
Thus, there is some subsequence, which we will just call $\Delta^{\scriptstyle{min}}_{\beta}$ for simplicity, and a (not yet known to be linear) function $\Delta$ such that $\Delta^{\scriptstyle{min}}_{\beta}\Lambda_{n}\xrightarrow{d}\Delta\Lambda$ and $(L\circ\Delta)\Lambda_{n}=\Lambda_{n}$ (Lemma 11.3) for each $n$. By $\|\Delta^{\scriptstyle{min}}_{\beta}\|\leq M$ and Lemma 7.1(c), $\Delta^{\scriptstyle{min}}_{\beta}\xrightarrow{SOT}\Delta$ and $\Delta$ is linear with $\|\Delta\|\leq M$. Finally, by boundedness of $L$, $L\circ\Delta=\operatorname{\mathbbm{1}}_{\mathsf{HP}(\mathcal{H})}$.
That $\Delta$ maps $\mathsf{cp}^{+}(\mathcal{H})$ into $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ follows just as in Lemma 11.5. ∎
11.4 The tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ and tangent space $\mathsf{cp}(\mathcal{H})$ are $d$-closed
In §10.2, we showed that $\mathsf{CP}(\mathcal{H})$ is $d$-closed. Now we will show the same holds for $\mathsf{cp}^{+}(\mathcal{H})$, and therefore for $\mathsf{cp}(\mathcal{H})$ as well, since $\mathsf{cp}(\mathcal{H})=\mathsf{cp}^{+}(\mathcal{H})\cap(-\mathsf{cp}^{+}(\mathcal{H}))$.
To that end, suppose $(\Lambda_{n})$ is a sequence in $\mathsf{cp}^{+}(\mathcal{H})$, $d$-convergent to, say, $\Lambda$. In addition, assume that this sequence is bounded in norm. Later, we will see how to remove this restriction. By Prop. 11.6, or even Lemma 11.5, there is a bounded sequence $Z_{n}$ of Lindblad parametrizations in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. This latter sequence has $d$-cluster points; let $Z$ be one such. Necessarily, it too is in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. By Lemma 11.4, $LZ$ is a $d$-cluster point of $(\Lambda_{n})$. Since $\Lambda_{n}\xrightarrow{d}\Lambda$, it follows that $LZ=\Lambda$.
Returning to the deferred business, drop the assumption that $(\Lambda_{n})$ is bounded. We show how to construct a bounded sequence also converging to $\Lambda$. By definition of $d$-convergence, there is $N_{1}$ such that $n>N_{1}$ implies that $\|\mathscr{P}_{1}(\Lambda-\Lambda_{n})\|<1$, $N_{2}>N_{1}$ such that $n>N_{2}$ implies that $\|\mathscr{P}_{2}(\Lambda-\Lambda_{n})\|<1$, etc. Now, define $\Gamma_{n}=\mathscr{P}_{n}\Lambda_{N_{n}}$. The sequence $(\Gamma_{n})$ is bounded, and still $d$-converges to $\Lambda$.
12 Conclusion
We have shown that the Jamiołkowski transform $\mathfrak{J}$ provides an excellent foundation on which to build the basic finite-dimensional theory of CP maps and of generators of Markovian CP evolutions in a unified way, and that the basic results can be extended to the separable setting by use of finite-dimensional approximations in the form of filtrations and their associated metrics. Among these basic results are: extreme CP maps are $\bm{\theta}$}-maps, every CP map has Kraus decompositions, a time-dependent element of $\mathsf{cp}^{+}(\mathcal{H})$ which is norm continuous generates a Markovian CP evolution, and Lindblad parametrizers exist. In addition, the classes $\mathsf{CP}(\mathcal{H},\mathcal{K})$, $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ are $d$-closed. It is significant that this extension can be done for Kraus decompositions, in particular, because there is an enormous gulf between the conventional approaches to the finite-dimensional and separable cases. Perhaps not so much as regards the GKSL problem, since Lindblad [2] does, in fact, make use of a sequence of finite-dimensional approximations in his key Proposition 5. However, there seems to be considerable difference in the way that primitive idea is used there compared to this paper.
acknowledgements
This work was supported by NSF MRSEC DMR-2011839. The author is grateful to the Open Quantum Systems discussion group at Penn State for stimulation.
Appendix A Glossary/Index
$\mathfrak{J}$ Jamiołkowski transform, §2.3.2
$\square\,$ $A\,\square\,B=\rho\mapsto A\rho B$, Def. 2.1
$\bm{\pi}$} ${\bm{\pi}}(h)=h\overline{h}$, Def. 2.2
$\bm{\theta}$} ${\bm{\theta}}(A)=A\,\square\,A^{\dagger}$, Def. 3.1
$\mathrm{SA}(\mathcal{H})$ hermitian operators, §3.1.1
$\mathrm{Pos}(\mathcal{H})$ positive operators, §3.1.2
$\mathsf{HP}(\mathcal{H})$ hermiticity-preserving superoperators, §3.1.3
$\mathsf{Mon}(\mathcal{H})$ monotone superoperators, §3.1.3
$\mathsf{CP}$ completely positive, Def. 3.4
(forward) tangent Def. 5.1
$\mathsf{cp}^{+}(\mathcal{H})$, $\mathsf{cp}(\mathcal{H})$ tangent cone/space to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$, Def. 5.2
$\Delta^{\scriptstyle{min}}$ Ex. 5.1
$L$-ops, $L$-parametrization, $L$-cone §5.3
$\mathsf{Lind\,{\mathcal{H}}}$ $\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$, (5.22)
Lindblad parametrizer Def. 5.3
$\mathcal{B}^{1}(\mathcal{H})$ trace-class operators, §7.1
SOT strong-operator topology, §7.2
WOT weak-operator topology, §7.3
filtration Def. 8.3
$\mathscr{P}$ projection in a (pre)filtration, Def. 8.1
$d$-metric metric induced by a filtration, §9
[…]
$\mathfrak{B}(\cdot,\cdot)$ bounded linear operators with operator norm
$\colon X xx$ belongs to XX”, as hypothesis
$(x_{n})$ sequence, implicitly indexed by $n$
$(x_{n})_{n}$ sequence, explicitly indexed by $n$
$\tau/\tau^{\prime}$ continuous continuous with topology $\tau$ ($\tau^{\prime}$) on domain (codomain)
$\mathrm{ball}(\mathscr{X})$ $\left\{x\colon\mathscr{X}\,\colon\,\|x\|<1\right\}$
$\overline{\mathrm{ball}}(\mathscr{X})$ $\left\{x\colon\mathscr{X}\,\colon\,\|x\|\leq 1\right\}$